### v1

# Mamba-FSCIL: Dynamic Adaptation with Selective State Space Model for Few-Shot Class-Incremental Learning

Abstract
Few-shot class-incremental learning (FSCIL) confronts the challenge of integrating new classes into a model with minimal training samples while preserving the knowledge of previously learned classes. Traditional methods widely adopt static adaptation relying on a fixed parameter space to learn from data that arrive sequentially, prone to overfitting to the current session. Existing dynamic strategies require the expansion of the parameter space continually, leading to increased complexity. To address these challenges, we integrate the recently proposed selective state space model (SSM) into FSCIL. Concretely, we propose a dual selective SSM projector that dynamically adjusts the projection parameters based on the intermediate features for dynamic adaptation. The dual design enables the model to maintain the robust features of base classes, while adaptively learning distinctive feature shifts for novel classes. Additionally, we develop a class-sensitive selective scan mechanism to guide dynamic adaptation. It minimizes the disruption to base-class representations caused by training on novel data, and meanwhile, forces the selective scan to perform in distinct patterns between base and novel classes. Experiments on miniImageNet, CUB-200, and CIFAR-100 demonstrate that our framework outperforms the existing state-of-the-art methods. The code is available at https://github.com/xiaojieli0903/Mamba-FSCIL.
1 Introduction
Figure 1: A sketch of our motivation. (a) Static adaptation with MLP is inadequate to separate the features belonging to different sessions that arrive incrementally, prone to overfitting to the current session classes; (b) Our method adopts selective SSM such that the projection learns an adaptive feature shift dependent on the input. Different color refers to data from different sessions.
Deep neural networks often struggle with adapting to new scenarios from limited data without forgetting previous knowledge. This is particularly challenging in real-time environments like robotics and autonomous driving, where data arrives incrementally and new categories often have few samples to learn. To address this challenge, the concept of few-shot class-incremental learning (FSCIL) was introduced [57, 76, 80]. In FSCIL, a model is trained on a base session with adequate labels and samples. Subsequently, it must adapt to new classes encountered in later sessions, each featuring only a few examples, without access to the data from previous sessions. Critically, the model needs to accurately classify images from all encountered classes in each session’s evaluation, balancing the adaptation to new classes with retaining old knowledge.
Various methods have been developed to address FSCIL. Data-based methods [29, 1, 42] replay data subsets or create synthetic samples to reinforce memory and mitigate catastrophic forgetting. Optimization-based techniques [7, 80, 56, 23, 33, 22, 5, 3, 10, 74, 9, 81] utilize advanced loss functions, regularization, and pseudo-feature generation to separate old-class and new-class features. However, most of these methods rely on static adaptation, which uses a pre-defined network architecture that does not change dynamically in response to new information. For example, as shown in Figure 1-(a), multi-layer perceptrons (MLPs) are usually used to project features from the frozen backbone to prototype classifiers [22, 42, 70]. While computationally efficient, MLP cannot dynamically adjust its parameters as class distributions evolve over different learning sessions. As a result, it has insufficient capacity to manage the delicate balance between preserving old knowledge (stability) and accommodating new ability (plasticity), leading to inferior performance on new classes or catastrophic forgetting of old ones. Dynamic network-based methods [57, 67, 66, 68, 76] offer an alternative by expanding architectures with each incremental session. They introduce extra task-specific parameters while preserving the parameters related to the old classes. While enhancing the capacity to incorporate new knowledge, these methods often require continuous structural changes and parameter expansion, leading to increased complexity and resource demands in incremental training.
Recently, Selective State Space Models (SSMs) [17], as known as the Mamba model, have emerged as a compelling alternative to the self-attention mechanisms [58, 11, 32] for efficiently handling long sequences. Mamba introduces a selective scan mechanism that dynamically computes its system matrices based on the input context, allowing for the effective handling of diverse and evolving data characteristics. This dynamic adaptation enables Mamba to adjust its operation parameters according to the needs of different input distributions, increasing the compatibility between old and new classes. Moreover, the 2D Selective Scan (SS2D) proposed for vision tasks [31, 55, 6, 47, 30] further enhances Mamba by spatially arranging image patches in four directions to ensure that each element of the feature map integrates information from all the other locations in different directions, thus creating a global receptive field without increasing linear computational complexity.
Inspired by the appealing merits of SSMs, this study explores how to fulfill the potential of Mamba to facilitate FSCIL. Unlike traditional FSCIL methods that rely on static structures or continually expanding parameters, Mamba is able to learn adaptive representation across old and new classes with a consistent parameter space dependent on the input. As shown in Figure 1-(b), by employing a Selective SSM-based projection layer that dynamically adjusts its parameters based on the input characteristics, the model can maintain the discriminative features of base classes, while compatibly integrating new classes without the need of expanding the parameter space. This not only preserves model efficiency but also minimizes the risk of overfitting, as the model does not accumulate excessive parameters that specialize narrowly on the training data of a particular session. Additionally, Vision Mamba’s [31, 83, 25, 65, 37] proficiency in processing visual tokens can capture fine-grained details of images, which is particularly beneficial when novel classes only have limited training samples, thereby enhancing the model’s perception of incremental classes through selective focus on the relevant image patches.
To this end, we propose Mamba-FSCIL in this paper, composed of a backbone network, a dual selective SSM projector, and a classifier fixed as a specific structure following [70]. The backbone network is used to learn a powerful feature extractor from the extensive training data in the base session and is frozen in the incremental sessions. Considering the huge disparity between the training sample quantities of the base session and incremental sessions, we develop a dual selective SSM projector to separately exploit the dynamic adaptation. Concretely, it consists of two projection branches and an identity branch. Each projection branch splits feature maps into patches and goes through an SS2D module after position embedding. The first one is used to project the backbone features into the classifier prototypes for base classes and is frozen in the incremental sessions. The other one is introduced since the first incremental session, to learn feature shifts of novel classes into their corresponding classifier prototypes.
To further enhance the dynamic adaptation effect, we propose a class-sensitive selective scan mechanism, which guides the selection pattern to be distinctive between the base classes and incremental classes. First, we use a suppression loss that regularizes the norm of gating for the features from the base-session classes. By doing so, the output feature of the base classes in incremental sessions will be close to the one in the base session, preserving the learned knowledge of the base classes as much as possible to induce stability. Second, we introduce a separation loss that minimizes the similarity of the system matrices (${\mathbf{B}}$, ${\mathbf{C}}$, and $\bm{\Delta}$) between the ones output from the base-class features and the ones output from all the incremental-class features. In this way, the selection in SS2D will be performed in different patterns for the base classes and the incremental classes, leaving sufficient feature space for the new classes in the incremental training process to induce plasticity.
To the best of our knowledge, we are the first to explore the potential of Mamba in the context of FSCIL. The contributions of this paper can be summarized as follows:
  • We successfully exploit the potential of Mamba for FSCIL by proposing a dual selective SSM projector. The dual design maintains the base-class features and adaptively adjusts its projection parameters based on the input context for dynamic adaptation on novel classes, mitigating catastrophic forgetting and enhancing adaptability.
  • We further develop a class-sensitive selective scan mechanism to guide the dynamic adaptation composed of two loss terms. The suppression loss minimizes the disruption to old-class representations caused by incremental training, maintaining stability for base classes. The separation loss makes the selective scan perform in distinct patterns to induce plasticity for novel classes.
  • Our approach is thoroughly evaluated on standard FSCIL benchmarks including CIFAR-100, CUB-200, and miniImageNet. Our results achieve state-of-the-art performances, demonstrating effectiveness in improving and preserving base-class accuracy and enhancing the adaptability to novel classes.
2 Related Works
2.1 Few-shot Class-Incremental Learning
FSCIL combines aspects of few-shot learning [45, 59], which focuses on learning from a small number of samples, with class-incremental learning [4, 26, 46], which involves adding new classes to a model’s knowledge base over time without retraining from scratch. FSCIL faces the key challenges: (1) Catastrophic Forgetting. Learning new classes without accessing the data of old classes can cause the model to forget previously acquired knowledge  [38, 15, 72]. (2) Data Scarcity. Limited data availability hampers the model’s ability to learn robust and generalizable representations and increases the overfitting risk [51, 54]. (3) Stability-Plasticity Dilemma. The model must seek a delicate balance, maintaining the knowledge of old classes while adapting to the new ones.
Various methods have been proposed to address the challenges. Data-based strategies mitigate catastrophic forgetting by replaying a subset of real data or generating synthetic data [29, 1, 42]. Optimization-based methods adopt techniques like meta-learning [74, 9, 81], pseudo-feature generation [8, 80], advanced loss functions or regularizers [56, 23, 33, 5, 3, 52, 2, 84, 16], well-designed classifiers [22, 70, 71, 82, 62], and knowledge distillation [57, 10, 7, 78]. For example, C-FSCIL [22] employs a mapping strategy that transforms input images into quasi-orthogonal prototypes to enhance class separability. NC-FSCIL [70] introduces a fixed equiangular tight frame classifier and trains a projection layer to attract the backbone features to these ideal prototypes, thus mitigating inter-class interference. However, most of these methods are based on static adaptation, whose parameters are not aware of the input, increasing the difficulty of striving for a comprehensive ability in both old and new classes. Existing dynamic adaptation methods dynamically adjust the model structure or interrelationships between prototypes [57, 67, 66, 68, 76]. TOPIC [57] uses a neural gas network to integrate new nodes and edges for new classes. CEC [76] employs graph models to evolve the classifier’s topology as new classes emerge. Dynamic networks such as LEC-Net [67] DER [66], and DSN [68] introduce task-specific parameters for new tasks while preserving those related to the old tasks.
Different from these approaches that often require expanding the parameter space, our method achieves dynamic adaptation leveraging the Selective SSM mechanism in Mamba that adjusts parameters based on input characteristics without increasing model complexity.
2.2 State Space Models
Transformers [32, 11, 21, 12], despite their proficiency in handling long-range dependencies, are hindered by their quadratic computational cost with respect to input size, particularly for long sequences or high-resolution images. In contrast, State Space Models (SSMs) provide an efficient alternative, capable of processing extensive sequences with significantly reduced computational demands [18, 19, 53, 44, 50, 41, 13, 61, 39, 63, 34]. SSMs involve recurrent updates over a sequence through hidden states, maintaining context in hidden states, and updating outputs by integrating these states with incoming inputs. Structured state space models (S4) [18, 19] exemplify this efficiency with a parameterized linear recurrent neural network. Subsequent research further enhances computational efficiency and expressive capacity of SSMs across various applications [19, 53, 44, 50, 41].
The recently proposed Mamba [17] enhances S4 with a data-dependent selective mechanism to selectively propagate or forget information according to the sequential input tokens and proposes a hardware-aware algorithm to boost efficiency. Mamba has also been adopted to process non-sequential input such as images and videos in vision tasks  [31, 83, 47, 30, 65, 37, 25]. For example, VMamba [31] introduces a cross-scan mechanism that processes images in both horizontal and vertical dimensions, allowing feature map elements to integrate information from all directions and providing a global receptive field without increasing linear computational complexity.
Although the Mamba model has been extensively applied in various domains such as text [17, 13, 18], vision [83, 31, 25, 35], audio [14], its potential to mitigate model forgetting has not yet been explored. This study is the first attempt to apply Mamba in FSCIL, effectively harnessing its capabilities of dynamic adaptation and sequence modeling to achieve state-of-the-art performance in FSCIL.
3 Preliminaries
3.1 Few-Shot Class-Incremental Learning
FSCIL trains a model incrementally in multiple sessions $\{{\mathcal{D}}^{(0)},{\mathcal{D}}^{(1)},\dots,{\mathcal{D}}^{(T)}\}$, where ${\mathcal{D}}^{(t)}=\{({\bm{x}}_{i},y_{i})\}_{i=1}^{|{\mathcal{D}}^{(t)}|}$ represents the training set for session $t$. ${\mathcal{D}}^{(0)}$ is the base session, and $T$ is the number of incremental sessions. The base session ${\mathcal{D}}^{(0)}$ usually contains a large label space ${\mathcal{C}}^{(0)}$ and extensive training data for each class $c\in{\mathcal{C}}^{(0)}$. In each incremental session ${\mathcal{D}}^{(t)}$, $t>0$, there are only a few labeled images, $|{\mathcal{D}}^{(t)}|=pq$, where $p$ is the number of classes and $q$ is the number of samples per novel class, known as $p$-way $q$-shot. There is no overlap in the label space between sessions i.e., ${\mathcal{C}}^{(t)}\cap{\mathcal{C}}^{(t^{\prime})}=\emptyset$ for all $t^{\prime}\neq t$. The training sets from previous sessions are not accessible, which requires the model to generalize to new data without forgetting previously learned information. Evaluation in session $t$ involves test data from all classes encountered up to that session, i.e., the label space of $\cup_{i=0}^{t}{\mathcal{C}}^{(i)}$.
3.2 Mamba
State Space Models (SSMs) map the input sequence $x(t)\in\mathbb{R}^{L}$ to the output sequences $y(t)\in\mathbb{R}^{L}$ through a recurrently updated hidden state $h(t)\in\mathbb{R}^{D}$, formulated as:
$h^{\prime}(t)={\mathbf{A}}h(t)+{\mathbf{B}}x(t),\quad y(t)={\mathbf{C}}h(t),$ (1)
where ${\mathbf{A}}\in\mathbb{R}^{D\times D}$ is the system’s evolution matrix, and ${\mathbf{B}}\in\mathbb{R}^{D\times 1}$, ${\mathbf{C}}\in\mathbb{R}^{1\times D}$ are the projection matrix. To adapt this continuous system for practical implementation, the structured state space model (S4) [18] and Mamba [17] employ the discretization transformation by zero-order hold rule, using a timescale parameter $\mathbf{\Delta}\in\mathbb{R}>0$ to transform the continuous parameters to discrete counterparts:
$\overline{{\mathbf{A}}}$ $=\exp{(\mathbf{\Delta}{\mathbf{A}})},$ (2)
$\overline{{\mathbf{B}}}$ $=(\mathbf{\Delta}{\mathbf{A}})^{-1}(\exp{(\mathbf{\Delta}{\mathbf{A}})}-\mathbf{I})\cdot\mathbf{\Delta}{\mathbf{B}}.$
The discretized system is defined as:
$h_{t}=\overline{{\mathbf{A}}}h_{t-1}+\overline{{\mathbf{B}}}x_{t},\quad y_{t}={\mathbf{C}}h_{t}.$ (3)
This system enables efficient global convolution operations that can be processed in parallel:
$\overline{{\mathbf{K}}}$ $=({\mathbf{C}}\overline{{\mathbf{B}}},{\mathbf{C}}\overline{{\mathbf{A}}}\overline{{\mathbf{B}}},\dots,{\mathbf{C}}\overline{{\mathbf{A}}}^{L-1}\overline{{\mathbf{B}}}),\quad\mathbf{y}={\mathbf{X}}*\overline{{\mathbf{K}}},$ (4)
where * denotes convolution operation, and $\overline{{\mathbf{K}}}\in\mathbb{R}^{L}$ is a convolutional kernel that covers the entire input sequence with the length of $L$. However, the current system employs constant operation parameters, e.g., ${{\mathbf{C}}}$ and ${{\mathbf{B}}}$, which remain static for varying input. This mechanism prevents S4 from selecting the correct information from the context or affecting the hidden state in an input-dependent manner.
Selective State Space Models (S6)
are the core component of Mamba, enhancing S4 by introducing a selective and dynamic mechanism for the interactions between sequential states [17]. Unlike conventional SSMs that utilize constant operation matrices that are agnostic of input features, S6 models utilize input-dependent parameters that calculate parameters ${\mathbf{B}}$, ${\mathbf{C}}$, and $\mathbf{\Delta}$ directly from the input sequence ${\bm{x}}$ through linear projections, enabling a dynamic and context-aware adaptation to the unique characteristics of the data.
2D Selective Scan (SS2D)
enhances Mamba’s ability to effectively capture spatial information for vision tasks [31]. SS2D arranges the feature patches in four different directions (top-left to bottom-right, bottom-right to top-left, top-right to bottom-left, and bottom-left to top-right), creating four separate sequences. Each sequence is processed by the S6 module before merging to reconstruct the 2D feature map. Given the input feature ${\bm{z}}$, the output feature $\overline{{\bm{z}}}$ of SS2D is formulated as:
$\overline{{\bm{z}}}=\text{SS2D}({\bm{z}})=\sum_{i=1}^{4}\text{S6}(\text{scan}({\bm{z}},i)),$ (5)
where $\text{scan}({\bm{z}},i)$ converts feature patches into sequence according to direction $i$. This approach ensures the comprehensive integration of contextual information from all directions, enhancing the model’s global receptive field.
4 Mamba-FSCIL
Figure 2: Mamba-FSCIL framework. The macro architecture (a) is composed of a backbone network, a dual selective SSM projector consisting of an identity branch and two selective SSM branches with the identical structure of (b), and a classifier. Our class-sensitive selective scan (c) performs $\mathcal{L}_{\text{supp}}$ and $\mathcal{L}_{\text{sep}}$ in the selective SSM branch to guide the dynamic adaptation in incremental session’s training.
4.1 Macro architecture
The architecture of Mamba-FSCIL is composed of a backbone network $f$, a dual selective SSM projector, and a linear classifier. The backbone network $f$ processes the input ${\bm{x}}$ to generate intermediate features ${\mathbf{F}}=f({\bm{x}})\in\mathbb{R}^{N\times D\times H\times W}$, where $N$, $D$, $H$, and $W$ denote the batch size, number of channels, height, and width, respectively. It is only trainable in the base session to learn a powerful feature extractor using the extensive training data of the base session. We design a dual selective SSM projector that maps the frozen intermediate features into the final representation for dynamic adaptation. The classifier adopts a fixed simplex equiangular tight frame (ETF) following [70, 79, 27, 49, 64, 36]. It provides an ideal feature space arrangement with the maximal equiangular separation among classes [69, 40].
4.2 Dual Selective SSM Projector
As illustrated in Fig. 2, a backbone network is trained in the base session to produce the intermediate feature ${\mathbf{F}}$, and will be frozen in the training of incremental sessions. To fully exploit the potential for dynamic adaptation, we develop a dual selective SSM projector consisting of three branches, i.e., an identity branch, and two selective SSM branches. The identity branch processes ${\mathbf{F}}$ through ${\bm{\mu}}^{\text{iden}}=p(\text{AvgPool}({\mathbf{F}}))$, where AvgPool is the average pooling function, $p$ is linear projection operation. In the selective SSM branch for the base session ($g^{\text{base}}$), ${\mathbf{F}}$ is first reshaped from a 2D feature map into a 1D vector to be processed by an MLP layer and is added with the position embedding as:
$\hat{{\mathbf{F}}}^{\text{base}}=\rm{MLP}(\text{Reshape}({\mathbf{F}}))+{\mathbf{E}}^{\text{base}},$ (6)
where ${\mathbf{E}}^{\text{base}}\in\mathbb{R}^{L\times D^{\prime}}$ denotes the position embedding, $\hat{{\mathbf{F}}}^{\text{base}}\in\mathbb{R}^{N\times L\times D^{\prime}}$ is the output feature, $D^{\prime}$ denotes the output dimension, and $L=H\times W$ is the spatial locations. Then $\hat{{\mathbf{F}}}^{\text{base}}$ goes through a selective SSM projector, which splits the features into two information streams via linear projections $f_{x}$ and $f_{z}$, creating ${\mathbf{X}}^{\text{base}}=f_{x}(\hat{{\mathbf{F}}}^{\text{base}})$ and ${\mathbf{Z}}^{\text{base}}=f_{z}(\hat{{\mathbf{F}}}^{\text{base}})$. ${\mathbf{X}}^{\text{base}}$ is intended for scanning, while ${\mathbf{Z}}^{\text{base}}$ is used for gating the output of the scanned results. Specifically, the ${\mathbf{X}}^{\text{base}}$ stream is then enhanced through a group convolution function followed by a $\rm{SiLU}$ activation function:
$\hat{{\mathbf{X}}}^{\text{base}}=\rm{SiLU}(\text{Conv}({\mathbf{X}}^{\text{base}})).$ (7)
The enhanced features $\hat{{\mathbf{X}}}^{\text{base}}$ are then processed by the SS2D module defined in Eq. 5, enhancing discrimination and stability crucial for maintaining high accuracy within base classes. The output of the selective SSM branch ${\bm{\mu}}^{\text{base}}$ is obtained by the element-wise multiplication of the SS2D module’s output and ${\mathbf{Z}}^{\text{base}}$ (after applying the $\rm{SiLU}$ activation) with average pooling:
${\bm{\mu}}^{\text{base}}=\text{AvgPool}\left(\text{SS2D}(\hat{{\mathbf{X}}}^{\text{base}})\odot\text{SiLU}({\mathbf{Z}}^{\text{base}})\right),$ (8)
where $\odot$ denotes element-wise multiplication. The SS2D system matrices for the input features $\hat{{\mathbf{X}}}^{\text{base}}$ are calculated by linear projections as:
${\mathbf{B}}=f_{B}(\hat{{\mathbf{X}}}^{\text{base}}),\quad{\mathbf{C}}=f_{C}(\hat{{\mathbf{X}}}^{\text{base}}),\quad\bm{\Delta}=f_{\Delta}(\hat{{\mathbf{X}}}^{\text{base}}),$ (9)
where ${\mathbf{B}}\in\mathbb{R}^{N\times K\times D_{B}\times L}$, ${\mathbf{C}}\in\mathbb{R}^{N\times K\times D_{C}\times L}$, $\bm{\Delta}\in\mathbb{R}^{N\times K\times D_{\Delta}\times L}$ are data-dependent system matrices, with $K$ representing the number of scan directions in Eq. 5. Here, $D_{*}$ represents the dimensions of each parameter. Then, the final representation of the base session is obtained by combining the outputs of the identity branch and the selective SSM branch as:
${\bm{\mu}}^{(0)}={\bm{\mu}}^{\text{base}}+{\bm{\mu}}^{\text{iden}}.$ (10)
During the base session, the backbone network and the selective SSM branch $g^{\text{base}}$ use the abundant available data to learn robust feature representations, thus laying a powerful feature extractor for later incremental training. Due to the huge gap in training data between the base session and incremental sessions, a single selective SSM is difficult to achieve a universal dynamic adaptation for all the base and incremental classes. To this end, our dual selective SSM projector adopts another selective SSM branch ($g^{\text{inc}}$) with an identical structure for incremental sessions.
Concretely, in each incremental session $1\leq t\leq T$, the backbone network $f$, the identity branch $p$, and the selective SSM branch for the base session $g^{\text{base}}$ are frozen, and only the newly introduced $g^{\text{inc}}$ is learnable to adapt the model to novel classes. Following the widely adopted practice in FSCIL [7, 5, 3, 22, 70], we employ a memory module ${\mathcal{M}}^{(t)}$, which stores the mean intermediate feature ${\mathbf{F}}_{c}$ of each old class $c$ to alleviate forgetting when encountering novel classes:
${\mathcal{M}}^{(t)}=\{{\mathbf{F}}_{c}\mid c\in\cup_{j=0}^{t-1}{\mathcal{C}}^{(j)}\},$ (11)
$\quad{\mathbf{F}}_{c}=\mathrm{Avg}_{i}\{f({\bm{x}}_{i})\mid y_{i}=c\}.$
The intermediate features of the new class data, together with the stored mean features of the old classes, go through $g^{\text{inc}}$ to obtain the output:
${\bm{\mu}}^{\text{inc}}=\text{AvgPool}\left(\text{SS2D}(\hat{{\mathbf{X}}}^{\text{inc}})\odot\text{SiLU}({\mathbf{Z}}^{\text{inc}})\right),$ (12)
where ${\mathbf{X}}^{\text{inc}}$ and ${\mathbf{Z}}^{\text{inc}}$ are processed with the same architecture as $g^{\text{base}}$. The resulting final representation for incremental sessions is aggregated as:
${\bm{\mu}}^{\text{(t)}}={\bm{\mu}}^{\text{base}}+{\bm{\mu}}^{\text{inc}}+{\bm{\mu}}^{\text{iden}},\quad t\geq 1.$ (13)
In this way, the branch $g^{\text{inc}}$ learns the shifts of feature distribution prompted by the new class data upon the robust representation established during the base session to dynamically accommodate the model to new classes. It enhances the adaptation ability of the model without altering the previously learned knowledge badly under the guidance of our class-sensitive selective scan.
4.3 Class-sensitive Selective Scan
In this section, we develop a class-sensitive selective scan mechanism that guides the dynamic adaptation of the selective SSM branch via two loss terms, i.e., the suppression loss and the separation loss.
Suppression Loss.
The feature vector ${\mathbf{Z}}^{\text{inc}}$ in  Eq. 12 plays a pivotal role in the element-wise multiplication with outputs from the SS2D module. By adjusting the norm of ${\mathbf{Z}}^{\text{inc}}$, we intend to manipulate the magnitude of the resulting feature vector ${\bm{\mu}}^{\text{inc}}$, which serves as a class-sensitive gating mechanism that controls the impact of new features with respect to base classes and incremental classes differently in the incremental session training.
  • For base classes, we reduce the norm of ${\mathbf{Z}}^{\text{inc}}$ of base classes to decrease ${\bm{\mu}}^{\text{inc}}$’s magnitude, thus minimizing feature shifts caused by the new branch $g^{\text{inc}}$ to maintain the stability of the features of the base class.
  • For novel classes, conversely, we increase the norm of ${\mathbf{Z}}^{\text{inc}}$ to enhance ${\bm{\mu}}^{\text{inc}}$’s magnitude, facilitating the learning of feature shifts that help the model adapt to new classes.
Therefore, the suppression loss can be formulated as:
${\mathcal{L}}^{\text{base}}_{\text{supp}}$ $=\sum_{({\mathbf{F}}_{c},y_{c})\in{\mathcal{M}}^{(0)}}\|{\mathbf{Z}}^{\text{inc}}_{c}\|^{2},$ (14)
${\mathcal{L}}^{\text{novel}}_{\text{supp}}$ $=-\sum_{({\bm{x}}_{i},y_{i})\in{\mathcal{D}}^{(t)}}\|{\mathbf{Z}}^{\text{inc}}_{i}\|^{2}-\sum_{({\mathbf{F}}_{c},y_{c})\in{{\mathcal{M}}}^{(t)}\backslash{{\mathcal{M}}}^{(0)}}\|{\mathbf{Z}}^{\text{inc}}_{c}\|^{2},$
where ${{\mathcal{M}}}^{(t)}\backslash{{\mathcal{M}}}^{(0)}=\{{\mathbf{F}}_{c}\mid c\in\cup_{j=1}^{t-1}{\mathcal{C}}^{(j)}\}$, and we have:
${\mathcal{L}}_{\text{supp}}={\mathcal{L}}^{\text{base}}_{\text{supp}}+{\mathcal{L}}^{\text{novel}}_{\text{supp}}.$ (15)
In this way, the interference with base-class features brought by the newly added branch $g_{\text{inc}}$ is diminished by managing ${\mathbf{Z}}^{\text{inc}}$ strategically. In the ideal case, it is expected that the output of $g_{\text{inc}}$, i.e., ${\bm{\mu}}^{\text{inc}}$, could be zero such that the inference of base-class data after incremental session’s training (Eq. (13)) will be equivalent to the one after the base session’s training (Eq. (10)) to induce stability.
Separation Loss.
We further introduce a separation term that incurs class sensitivity in the input-dependent parameters (${\mathbf{B}},{\mathbf{C}},\bm{\Delta}$) in SS2D. For both base and novel classes, we calculate their average input-dependent parameters:
$\bar{{\mathbf{P}}}_{\text{b}}$ $=\frac{1}{N_{\text{base}}\cdot K\cdot L}\sum_{({\mathbf{F}}_{i},y_{i})\in{\mathcal{M}}^{(0)}}\sum_{j=1}^{K}\sum_{k=1}^{L}{\mathbf{P}}_{i,j,:,k},$ (16)
$\bar{{\mathbf{P}}}_{\text{n}}$ $=\frac{1}{N_{\text{novel}}\cdot K\cdot L}\sum_{{\mathcal{D}}^{(t)}\cup({{\mathcal{M}}}^{(t)}\backslash{{\mathcal{M}}}^{(0)})}\sum_{j=1}^{K}\sum_{k=1}^{L}{\mathbf{P}}_{i,j,:,k},$
where ${\mathbf{P}}$ represents any of ${\mathbf{B}}$, ${\mathbf{C}}$, or $\bm{\Delta}$ and $N_{\text{base}}$ and $N_{\text{novel}}$ denote the number of samples in base and novel classes, respectively. We then compute the cosine similarity of these average parameters between $\bar{{\mathbf{P}}}_{\text{b}}$ and $\bar{{\mathbf{P}}}_{\text{n}}$, aiming to minimize this similarity to maximize the separability of these selective scan parameters between base and novel classes. This process can be formulated as:
${\mathcal{L}}_{\text{sep}}=\frac{\bar{{\mathbf{B}}}_{\text{b}}\cdot\bar{{\mathbf{B}}}_{\text{n}}}{\|\bar{{\mathbf{B}}}_{\text{b}}\|\|\bar{{\mathbf{B}}}_{\text{n}}\|}+\frac{\bar{{\mathbf{C}}}_{\text{b}}\cdot\bar{{\mathbf{C}}}_{\text{n}}}{\|\bar{{\mathbf{C}}}_{\text{b}}\|\|\bar{{\mathbf{C}}}_{\text{n}}\|}+\frac{\bar{\bm{\Delta}}_{\text{b}}\cdot\bar{\bm{\Delta}}_{\text{n}}}{\|\bar{\bm{\Delta}}_{\text{b}}\|\|\bar{\bm{\Delta}}_{\text{n}}\|}.$ (17)
As a result, the selective scan can be performed in distinct patterns between base and novel classes to induce plasticity in the incremental session’s training.
4.4 Optimization
After we get the final representation, i.e., Eq. (10) for the base session and Eq. (13) for incremental sessions, we calculate the classification loss with the ETF classifier. Following [70, 49], we use the dot regression (DR) loss [69] as the training objective, as it is shown to outperform cross entropy loss when using a fixed ETF classifier in imbalanced training, which FSCIL naturally satisfies. The DR loss can be written as:
${\mathcal{L}}_{\text{DR}}\left(\hat{{\bm{\mu}}}_{i},\hat{{\mathbf{W}}}_{\rm ETF}\right)=\frac{1}{2}\left(\hat{\mathbf{w}}^{T}_{y_{i}}\hat{{\bm{\mu}}}_{i}-1\right)^{2},$ (18)
where $\hat{{\bm{\mu}}}_{i}=\frac{{\bm{\mu}}_{i}}{\|{\bm{\mu}}_{i}\|}$ is the $l_{2}$ normalized feature of the final representation, $y_{i}$ is the label of input $x_{i}$, and $\hat{\mathbf{w}}_{y_{i}}$ is the classifier prototype in $\hat{{\mathbf{W}}}_{\rm ETF}$ for label $y_{i}$.
Base Session Training.
During base session training ($t=0$), the backbone network $f$, the identity branch $p$, and the selective SSM branch for the base session $g^{\text{base}}$ are jointly optimized to establish a robust feature extractor. The objective during this phase is to minimize the classification loss on the extensive base session dataset:
$\min_{f,p,g^{\text{base}}}\frac{1}{|{\mathcal{D}}^{(0)}|}\sum_{({\bm{x}}_{i},y_{i})\in{\mathcal{D}}^{(0)}}{\mathcal{L}}_{\text{DR}}\left(\hat{{\bm{\mu}}}_{i}^{(0)},\hat{{\mathbf{W}}}_{\rm ETF}\right).$ (19)
Incremental Sessions Training.
In subsequent sessions ($1\leq t\leq T$), as new classes are introduced, the optimization of the model focuses on seamlessly integrating these classes while minimizing disruption of previously learned knowledge. The training objective during these sessions is to adapt the model with new-class data and maintain the existing ability of old classes with the stored intermediate features in ${\mathcal{M}}^{(t)}$. Thus, the classification loss can be formulated as:
${\mathcal{L}}_{\text{cls}}^{(t)}=\frac{1}{|{\mathcal{D}}^{(t)}|+|{\mathcal{M}}^{(t)}|}\left({\mathcal{L}}^{(t)}_{\text{cls-new}}+{\mathcal{L}}^{(t)}_{\text{cls-old}}\right),$ (20)
where
${\mathcal{L}}^{(t)}_{\text{cls-new}}$ $=\sum_{({\bm{x}}_{i},y_{i})\in{\mathcal{D}}^{(t)}}{\mathcal{L}}_{\text{DR}}\left(\hat{{\bm{\mu}}}_{i}^{(t)},\hat{{\mathbf{W}}}_{\rm ETF}\right),$ (21)
${\mathcal{L}}^{(t)}_{\text{cls-old}}$ $=\sum_{({\mathbf{F}}_{c},y_{c})\in{\mathcal{M}}^{(t)}}{\mathcal{L}}_{\text{DR}}\left(\hat{{\bm{\mu}}}_{c}^{(t)},\hat{{\mathbf{W}}}_{\rm ETF}\right).$
Together with the suppression loss and the separation loss in Section 4.3, the total objective in session $t\geq 1$ is:
$\min_{g^{\text{inc}}}\left({\mathcal{L}}_{\text{cls}}^{(t)}+\alpha{\mathcal{L}}_{\text{supp}}^{(t)}+\beta{\mathcal{L}}_{\text{sep}}^{(t)}\right),$ (22)
where $\alpha$ and $\beta$ are hyper-parameters to balance the loss scales.
5 Experiments
Our experiments evaluate Mamba-FSCIL on the three wildly used benchmarks including miniImageNet [48], CIFAR-100 [24], and CUB-200 [60], comparing it against static and dynamic methods. We conduct ablation studies to assess the impact of our dual selective SSM projector and class-sensitive selective scan mechanism. We also provide an analysis of feature embeddings and compare our dynamic adaptation ability with Dynamic Support Networks (DSN) [68] to demonstrate efficiency in model complexity and superior performances across sessions.
5.1 Datasets and Implementation Details
Datasets. We employ three benchmark datasets. miniImageNet is a variant of ImageNet with 100 classes, each having 500 training and 100 testing images of $84\times 84$ pixels. CIFAR-100 has the same number of classes and images, and the image size is $32\times 32$. CUB-200 is a fine-grained classification dataset consisting of 11,788 images in 200 classes, with an image resolution of $224\times 224$ pixels. We follow the standard experimental settings in FSCIL [57, 76, 70]. For both miniImageNet and CIFAR-100, the base session contains 60 classes, and a 5-way 5-shot problem (5 classes and 5 images per class) is adopted for each of the 8 incremental sessions. For CUB-200, 100 classes are used in the base session, and there are 10 incremental sessions, each of which is a 10-way 5-shot.
Training Details. We employ the standard data preprocessing and augmentation techniques, including random resizing, flipping, and color jittering [57, 76, 42, 70]. We conduct training with a batch size of 512 for base sessions and 64 for incremental sessions, which includes new session data along with intermediate features stored in memory. For miniImageNet, we utilize ResNet12 [20] as the backbone and train for 500 epochs in the base session, followed by 100-170 iterations in each incremental session. The initial learning rate is 0.25 for the base session and 0.01 for the incremental sessions. For CIFAR-100, we utilize ResNet12 [20] as the backbone and train the base session for 200 epochs, followed by 200 iterations per incremental session. The initial learning rate is 0.25 for both base and incremental sessions. For CUB-200, we utilize ResNet18 [20] as backbone and train for 80 epochs in the base session, followed by 200-290 iterations in each incremental session. The initial learning rates are 0.02 for the backbone and 0.2 for the projector during the base session, and 0.05 for each incremental session. We apply a cosine annealing strategy for the learning rate and use SGD with a momentum of 0.009 and a weight decay of 0.0005 as the optimizer. The hyper-parameters $\alpha$ and $\beta$ are set between $50$ and $200$, $0.05$ and $0.5$, respectively.
Table 1: Performance of FSCIL in each session on miniImageNet and comparison with other studies. “Average Acc.” is the average accuracy of all sessions. “Final Improv.” calculates the improvement of our method in the last session over previous methods.
Methods Accuracy in each session (%) $\uparrow$ Average Final
0 1 2 3 4 5 6 7 8 Acc. Improv.
Static adaptation methods
IDLVQ [5] 64.77 59.87 55.93 52.62 49.88 47.55 44.83 43.14 41.84 51.16 +17.52
Self-promoted [82] 61.45 63.80 59.53 55.53 52.50 49.60 46.69 43.79 41.92 52.76 +17.44
SFbFSCIL [8] 61.40 59.80 54.20 51.69 49.45 48.00 45.20 43.80 42.10 50.63 +17.26
Data-free Replay [29] 71.84 67.12 63.21 59.77 57.01 53.95 51.55 49.52 48.21 58.02 +11.15
LIMIT [81] 72.32 68.47 64.30 60.78 57.95 55.07 52.70 50.72 49.19 59.06 +10.17
MetaFSCIL [9] 72.04 67.94 63.77 60.29 57.58 55.16 52.90 50.79 49.19 58.85 +10.17
C-FSCIL [22] 76.40 71.14 66.46 63.29 60.42 57.46 54.78 53.11 51.41 61.61 +7.95
Regularizer [3] 80.37 74.68 69.39 65.51 62.38 59.03 56.36 53.95 51.73 63.71 +7.63
CABD [78] 74.65 70.43 66.29 62.77 60.75 57.24 54.79 53.65 52.22 61.42 +7.14
ALICE [42] 80.60 70.60 67.40 64.50 62.50 60.00 57.80 56.80 55.70 63.99 +3.66
SAVC [52] 81.12 76.14 72.43 68.92 66.48 62.95 59.92 58.39 57.11 67.05 +2.25
NC-FSCIL [70] 84.02 76.80 72.00 67.83 66.35 64.04 61.46 59.54 58.31 67.82 +1.05
Dynamic adaptation methods
TOPIC [57] 61.31 50.09 45.17 41.16 37.48 35.52 32.19 29.46 24.42 39.64 +34.94
DSN [68] 68.95 63.46 59.78 55.64 52.85 51.23 48.9 46.78 45.89 54.83 +13.47
CEC [76] 72.00 66.83 62.97 59.43 56.70 53.73 51.19 49.24 47.63 57.75 +11.73
Mamba-FSCIL (ours) 84.93 80.02 74.61 71.33 69.15 65.62 62.38 60.93 59.36 69.81
Table 2: Performance of FSCIL in each session on CIFAR-100 and comparison with other studies. “Average Acc.” is the average accuracy of all sessions. “Final Improv.” calculates the improvement of our method in the last session over previous methods.
Methods Accuracy in each session (%) $\uparrow$ Average Final
0 1 2 3 4 5 6 7 8 Acc. Improv.
Static adaptation methods
Self-promoted [82] 64.10 65.86 61.36 57.45 53.69 50.75 48.58 45.66 43.25 54.52 +14.26
Data-free Replay [29] 74.40 70.20 66.54 62.51 59.71 56.58 54.52 52.39 50.14 60.78 +7.37
MetaFSCIL [9] 74.50 70.10 66.84 62.77 59.48 56.52 54.36 52.56 49.97 60.79 +7.54
C-FSCIL [22] 77.47 72.40 67.47 63.25 59.84 56.95 54.42 52.47 50.47 61.64 +7.04
LIMIT [81] 73.81 72.09 67.87 63.89 60.70 57.77 55.67 53.52 51.23 61.84 +6.28
ALICE [42] 79.00 70.50 67.10 63.40 61.20 59.20 58.10 56.30 54.10 63.21 +3.41
CABD [78] 79.45 75.38 71.84 67.95 64.96 61.95 60.16 57.67 55.88 66.14 +1.63
NC-FSCIL [70] 82.52 76.82 73.34 69.68 66.19 62.85 60.96 59.02 56.11 67.50 +1.40
Dynamic adaptation methods
TOPIC [57] 64.10 55.88 47.07 45.16 40.11 36.38 33.96 31.55 29.37 42.62 +28.14
CEC [76] 73.07 68.88 65.26 61.19 58.09 55.57 53.22 51.34 49.14 59.53 +8.37
DSN [68] 73.00 68.83 64.82 62.24 59.16 56.96 54.04 51.57 49.35 60.00 +8.16
Mamba-FSCIL (ours) 82.89 77.85 73.69 69.67 66.89 63.66 61.48 59.74 57.51 68.12
Table 3: Performance of FSCIL in each session on CUB-200 and comparison with other studies.“Average Acc.” is the average accuracy of all sessions. “Final Improv.” calculates the improvement of our method in the last session over previous methods.
Methods Accuracy in each session (%) $\uparrow$ Average Final
0 1 2 3 4 5 6 7 8 9 10 Acc. Improv.
Static adaptation methods
Fixed classifier [43] 68.47 51.00 45.42 40.76 35.90 33.18 27.23 24.24 21.18 17.34 16.20 34.63 +45.45
Self-promoted [82] 68.68 61.85 57.43 52.68 50.19 46.88 44.65 43.07 40.17 39.63 37.33 49.32 +24.32
SFbFSCIL [8] 68.78 59.37 59.32 54.96 52.58 49.81 48.09 46.32 44.33 43.43 43.23 51.84 +18.42
DeepInv [73] 75.90 70.21 65.36 60.14 58.79 55.88 53.21 51.27 49.38 47.11 45.67 57.54 +15.98
D-DeepEMD [75] 75.35 70.69 66.68 62.34 59.76 56.54 54.61 52.52 50.73 49.20 47.60 58.73 +14.05
D-NegCosine [28] 74.96 70.57 66.62 61.32 60.09 56.06 55.03 52.78 51.50 50.08 48.47 58.86 +13.18
Data-free replay [29] 75.90 72.14 68.64 63.76 62.58 59.11 57.82 55.89 54.92 53.58 52.39 61.52 +9.26
MetaFSCIL [9] 75.9 72.41 68.78 64.78 62.96 59.99 58.3 56.85 54.78 53.82 52.64 61.93 +9.01
MgSvF [77] 72.29 70.53 67.00 64.92 62.67 61.89 59.63 59.15 57.73 55.92 54.33 62.37 +7.32
FACT [80] 75.90 73.23 70.84 66.13 65.56 62.15 61.74 59.83 58.41 57.89 56.94 64.42 +4.71
IDLVQ [5] 77.37 74.72 70.28 67.13 65.34 63.52 62.10 61.54 59.04 58.68 57.81 65.23 +3.84
LIMIT [81] 76.32 74.18 72.68 69.19 68.79 65.64 63.57 62.69 61.47 60.44 58.45 66.67 +3.20
NC-FSCIL [70] 80.45 75.98 72.30 70.28 68.17 65.16 64.43 63.25 60.66 60.01 59.44 67.28 +2.21
ALICE [42] 77.40 72.70 70.60 67.20 65.90 63.40 62.90 61.90 60.50 60.60 60.10 65.75 +1.55
Dynamic adaptation methods
TOPIC [57] 68.68 62.49 54.81 49.99 45.25 41.40 38.35 35.36 32.22 28.31 26.28 43.92 +35.37
CEC [76] 75.85 71.94 68.50 63.50 62.43 58.27 57.73 55.81 54.83 53.52 52.28 61.33 +9.37
DSN [68] 76.06 72.18 69.57 66.68 64.42 62.12 60.16 58.94 56.99 55.10 54.21 63.31 +7.44
Mamba-FSCIL (ours) 80.90 76.26 72.97 70.14 67.83 65.74 65.43 64.12 62.31 62.12 61.65 68.22
5.2 Comparison with State-of-the-art Methods
The comparison with the state-of-the-art methods on minImageNet, CIFAR-100, and CUB-200 is shown in Tab. 1, Tab. 2, and Tab. 3. The results demonstrate that Mamba-FSCIL consistently outperforms existing static and dynamic adaptation methods by over 1% in terms of the last-session accuracy, where all classes are taken into account. Specifically, on miniImageNet, Mamba-FSCIL achieves the highest average accuracy of 69.81% and increases the last-session accuracy to 59.36%, which surpasses the challenging baseline NC-FSCIL [70] by 1.99% and 1.05%, respectively. In CIFAR-100, Mamba-FSCIL maintains the best performance in most sessions, with the last session accuracy outperforming NC-FSCIL by 1.4%. Notably, Mamba-FSCIL significantly surpasses dynamic methods like DSN [68] with an average accuracy improvement of 8.12% over DSN. In the more challenging fine-grained CUB-200 dataset, Mamba-FSCIL leads with an average accuracy of 68. 22%, outperforming all static and dynamic methods. Compared to DSN, Mamba-FSCIL shows the remarkable ability to learn complex classification scenarios incrementally, improving the accuracy in the last session by 7.44%, also indicating a lower performance drop over sessions.
5.3 Ablation Studies
Table 4: Ablation studies to investigate the effects of our dual selective projector and the class-sensitive selective scan mechanism on FSCIL performance across different datasets. “Base” refers to the accuracy of the first session; “Final” refers to the last session; “PD” denotes the performance drop, i.e., the accuracy difference between the first and the last session.
Methods Adaptation Type miniImageNet CIFAR-100 CUB-200
[…]
Baseline Static 84.02 58.31 25.71 82.52 56.11 26.41 80.45 59.44 21.01
+ Single Selective SSM Projector Dynamic 84.93 57.89 27.04 82.80 56.07 26.73 80.90 60.70 20.20
+ Dual Selective SSM Projector Dynamic 84.93 58.92 26.01 82.80 56.71 26.09 80.90 61.06 19.84
+ Dual Selective SSM Projector + Class-sensitive Selective Scan Dynamic 84.93 59.36 25.57 82.80 57.51 25.29 80.90 61.65 19.25
In our ablation studies across the three datasets, we assess the impact of our dual selective projector and class-sensitive selective scan mechanism. Utilizing a consistent setup that includes the same backbone, identity branch, memory module, classifier, and classification loss, we explore several configurations: (1) Baseline employs NC-FSCIL with a static MLP-based projection layer, serving as a competitive static adaptation method. (2) + Single Selective SSM Projector integrates a single dynamic SSM projector with one projection branch, updated across all sessions. (3) + Dual Selective SSM Projector uses our dual selective SSM projector with two branches, as introduced in Sec. 4.2. (4) + Dual Selective SSM Projector + Class-sensitive Selective Scan further integrates our class-sensitive selective scan.
The introduction of a single selective SSM projector improves base session accuracy across the three datasets (e.g., from 84.02% to 84.93% on miniImageNet and from 82.52% to 82.80% on CIFAR-100), demonstrating the Mamba-based projector helps to improve the performance in the base session, providing a more powerful feature extractor for the subsequent sessions. However, despite these gains, the final performances on miniImageNet and CIFAR-100 both fall below the baseline, implying that a naive replacement of the MLP projector with the selective SSM projector cannot effectively mitigate catastrophic forgetting, increasing the performance drop (PD) as a result. In contrast, when our proposed dual selective SSM projector is adopted, the final performance is also effectively improved over the baseline, reducing the performance drop from 27.04% to 26.01% on miniImageNet and from 26.73% to 26.09% on CIFAR-100. This shows that our dual design is capable of adaptively learning feature shifts of novel classes into their corresponding classifier prototypes with fewer disruptions to the learned knowledge of the base classes. Further improvements are achieved with the integration of our class-sensitive scan mechanism into the dual projector, increasing the last-session accuracy from 56.71% to 57.51% on CIFAR-100 and from 61.06% to 61.65% on CUB-200. These findings validate the critical role of each component in Mamba-FSCIL on multiple datasets.
We also conduct a sensitivity analysis on the hyper-parameters $\alpha$ and $\beta$ of Eq. (22) in Tab. 5 to evaluate their impact on our model’s average accuracy on miniImageNet. Each experiment varies one parameter with the other fixed. The results show that when $\alpha$ and $\beta$ are set within a proper range, the performance only varies slightly, demonstrating the model’s robustness to the hyper-parameters.
Table 5: Effects of hyper-parameters $\alpha$ and $\beta$.
Impact of hyper-parameter $\alpha$
Value 40 80 120 140 160
Average Acc. 69.72 69.81 69.81 69.80 69.79
Impact of hyper-parameter $\beta$
Value 0.05 0.1 0.3 0.5 0.7
Average Acc. 69.64 69.70 69.71 69.81 69.79
5.4 Analysis of Feature Embeddings
We analyze feature embeddings using t-SNE visualizations for both base and novel classes from the test set of the CIFAR-100 dataset, comparing the strong baseline NC-FSCIL with our Mamba-FSCIL. Fig. 3-(a) and Fig. 3-(b) show the results of some base classes during the first and last sessions. It is shown that the base-class features of the baseline method are much scattered in the last session. As a comparison, Mamba-FSCIL demonstrates consistent intra-class compactness and distinct inter-class separation, indicating enhanced stability for base classes that better maintain their performance in incremental training. Fig. 3-(c) and Fig. 3-(d) depict the results of some novel classes in the 6-th and 8-th sessions. In both sessions, Mamba-FSCIL shows significantly more compact clusters with clearer margins among classes, highlighting its capacity to accommodating new classes. These results explain Mamba-FSCIL’s superior performance in the benchmarks.
5.5 Components in Class-sensitive Selective Scan
We investigate the impacts of the suppression and separation losses introduced in Sec. 4.3. First, we measure the average cosine similarity between the final representations and the classifier prototypes across all base classes in Fig. 4, defined as:
$\frac{1}{\sum_{k=1}^{K}N_{k}}\sum_{k=1}^{K}\sum_{i=1}^{N_{k}}\cos\angle({\bm{\mu}}_{k,i},\mathbf{w}_{k}),$ (23)
where ${\bm{\mu}}_{k,i}$ represents the feature vector of sample $i$ in class $k$, $N_{k}$ is the number of samples in class $k$, $\mathbf{w}_{k}$ is the classifier prototype for class $k$, and $K$ is the number of base classes. To assess the separability between base and novel classes, we calculate the average cosine similarity for any possible pairing of base and novel classes feature vectors as:
$\frac{1}{N_{\text{base}}N_{\text{novel}}}\sum_{i=1}^{N_{\text{base}}}\sum_{j=1}^{N_{\text{novel}}}\cos\angle({\bm{\mu}}_{\text{base},i},{\bm{\mu}}_{\text{novel},j}),$ (24)
where ${\bm{\mu}}_{\text{base},i}$ and ${\bm{\mu}}_{\text{novel},j}$ are feature vectors of the base class sample $i$ and novel class sample $j$, respectively.
Fig. 4 shows that the incorporation of suppression loss ${\mathcal{L}}_{\text{supp}}$ helps to improve the base-class alignment between features and classifier prototypes in late sessions, facilitating the maintenance of base-class performance. When the separation loss ${\mathcal{L}}_{\text{sep}}$ is further added, the similarities of features between the base and novel classes are significantly reduced, which leaves sufficient feature space for incremental training sessions, promoting the novel-class accuracy.
Figure 3: t-SNE visualization of feature embeddings from the CIFAR-100 dataset, showcasing base classes in (a) and (b) and novel classes in (c) and (d). Each color denotes a different class.
Figure 4: Impact of suppression and separation losses on the alignment between base-class feature and classifier and the separability between base and novel classes in MiniImageNet’s test set: (a) average cosine similarity of base class features with their corresponding classifiers; (b) average cosine similarity between base-class and novel-class features.
Figure 5: Comparison of parameter growth (a) and session-wise accuracy (b) between Mamba-FSCIL and DSN [68].
5.6 Comparison with Dynamic Adaptation Method
We compare our Mamba-FSCIL with DSN [68] in Fig. 5, particularly in terms of parameter efficiency and session-wise performance. Both methods use the same ResNet18 backbone. As shown in Fig. 5-(a), Unlike DSN, which continually increases its parameter amount to learn new classes, Mamba-FSCIL armed with a dual selective SSM projector dynamically adjusts its parameters without expanding the parameter space. In addition, Fig. 5-(b) illustrates that Mamba-FSCIL consistently outperforms DSN in all sessions, demonstrating Mamba-FSCIL’s superior ability to retain old knowledge while learning new classes.
6 Conclusion
In this study, we propose the Mamba-FSCIL framework, leveraging selective state space models to address the challenges of FSCIL. Mamba-FSCIL utilizes a dual selective SSM projector to achieve dynamic adaptation without the need to expand the model’s parameter space continually. The dual design dynamically adjusts its parameters to integrate new classes while preserving the integrity of previously learned information. Furthermore, the incorporation of the class-sensitive selective scan mechanism, composed of the suppression loss and the separation loss, guides the dynamic adaptation process deliberately, enhancing both stability and adaptability in incremental sessions’ training. Empirical results demonstrate that Mamba-FSCIL outperforms the state-of-the-art static and dynamic methods across three benchmark datasets.