### v2

# Mamba-FSCIL: Dynamic Adaptation with Selective State Space Model for Few-Shot Class-Incremental Learning

Abstract
Few-shot class-incremental learning (FSCIL) confronts the challenge of integrating new classes into a model with minimal training samples while preserving the knowledge of previously learned classes. Traditional methods widely adopt static adaptation relying on a fixed parameter space to learn from data that arrive sequentially, prone to overfitting to the current session. Existing dynamic strategies require the expansion of the parameter space continually, leading to increased complexity. In this study, we explore the potential of Selective State Space Models (SSMs) for FSCIL, leveraging its dynamic weights and strong ability in sequence modeling to address these challenges. Concretely, we propose a dual selective SSM projector that dynamically adjusts the projection parameters based on the intermediate features for dynamic adaptation. The dual design enables the model to maintain the robust features of base classes, while adaptively learning distinctive feature shifts for novel classes. Additionally, we develop a class-sensitive selective scan mechanism to guide dynamic adaptation. It minimizes the disruption to base-class representations caused by training on novel data, and meanwhile, forces the selective scan to perform in distinct patterns between base and novel classes. Experiments on miniImageNet, CUB-200, and CIFAR-100 demonstrate that our framework outperforms the existing state-of-the-art methods. The code is available at https://github.com/xiaojieli0903/Mamba-FSCIL.
Index Terms:
Few-Shot Class Incremental Learning, Selective State Space Model, Dynamic Adaptation, Class Incremental Learning
1 Introduction
Deep neural networks often struggle with adapting to new scenarios from limited data without forgetting previous knowledge. This is particularly challenging in real-time environments like robotics and autonomous driving, where data arrives incrementally and new categories often have few samples to learn. To address this challenge, the concept of few-shot class-incremental learning (FSCIL) was introduced [1, 2, 3]. In FSCIL, a model is trained on a base session with sufficient labels and samples. Subsequently, it must adapt to new classes encountered in later sessions, each containing only a few examples, without access to the data from previous sessions. Critically, the model needs to accurately classify images from all encountered classes in each session’s evaluation, balancing the adaptation to new classes with retaining knowledge of previously learned classes.
Various methods have been developed to address FSCIL [4, 5]. Data-based methods [6, 7, 8] replay data subsets or create synthetic samples to reinforce memory and mitigate catastrophic forgetting. Optimization-based techniques [9, 10, 11, 12, 13, 14, 15, 16, 17, 18] utilize advanced loss functions, regularization, and pseudo-feature generation to separate old-class and new-class features. However, most of these methods rely on static adaptation, which uses a pre-defined network architecture that does not change dynamically in response to new information. For example, as shown in Figure 1 (a), multi-layer perceptrons (MLPs) are usually used to project features from the frozen backbone to prototype classifiers [8, 16, 19]. While computationally efficient, MLP cannot dynamically adjust its parameters as class distributions evolve over different learning sessions. As a result, it has insufficient capacity to manage the delicate balance between preserving old knowledge (stability) and accommodating new ability (plasticity), leading to inferior performance on new classes or catastrophic forgetting of old ones. Dynamic network-based methods [1, 2, 20, 21, 22, 23, 24, 25, 26] offer an alternative by expanding architectures with each incremental session. They introduce extra task-specific parameters while preserving the parameters related to the old classes. While enhancing the capacity to incorporate new knowledge, these methods often require continuous structural changes and parameter expansion, leading to increased complexity and resource demands in incremental training.
Figure 1: A sketch of our motivation. (a) Static Adaptation with MLP: MLPs apply the same transformation to all inputs, which is inadequate for separating evolving features from different sessions and prone to overfitting to current session classes; (b) Dynamic Adaptation with Selective SSMs: Selective SSMs dynamically adjust parameters based on input $x_{t}$, generating input-dependent parameters ($\bm{\Delta}_{t}$, ${\mathbf{B}}_{t}$, ${\mathbf{C}}_{t}$) to learn adaptive feature shifts in the output $y_{t}$. Different colors (red, green, purple) represent data from different sessions.
Recently, Selective State Space Models (SSMs) [27], as known as the Mamba model, have emerged as a compelling alternative to self-attention mechanisms [28, 29, 30] to efficiently handle long sequences. Mamba introduces a selective scan mechanism that dynamically computes its system matrices based on the input context, allowing for the effective handling of diverse and evolving data characteristics. This dynamic adaptation enables Mamba to adjust its operation parameters according to the needs of different input distributions, increasing the compatibility between old and new classes. Moreover, the 2D Selective Scan (SS2D) proposed for vision tasks [31, 32, 33, 34] further enhances Mamba by spatially arranging image patches in four directions to ensure that each element of the feature map integrates information from all other locations in different directions, thus creating a global receptive field without increasing linear computational complexity.
Inspired by the appealing merits of SSMs, this study explores how to fulfill the potential of Mamba to facilitate FSCIL. Unlike traditional FSCIL methods that rely on static structures or continually expanding parameters, Mamba is able to learn adaptive representation across old and new classes with a consistent parameter space dependent on the input. As shown in Figure 1 (b), by employing a Selective SSM-based projection layer that dynamically adjusts its parameters based on input characteristics, the model can maintain the discriminative features of the base classes, while compatibly integrating new classes without the need of expanding the parameter space. This not only preserves model efficiency but also minimizes the risk of overfitting, as the model does not accumulate excessive parameters that specialize narrowly on the training data of a particular session. Additionally, Vision Mamba’s [31, 35, 36, 37, 38, 39] proficiency in processing visual tokens can capture fine-grained details of images, which is particularly beneficial when novel classes only have limited training samples, thereby enhancing the model’s perception of incremental classes through selective focus on the relevant image patches.
To this end, we propose Mamba-FSCIL in this paper, composed of a backbone network, a dual selective SSM projector, and a classifier fixed as a specific structure following [19]. The backbone network is used to learn a powerful feature extractor from the extensive training data in the base session and is frozen in the incremental sessions. Considering the huge disparity between the training sample quantities of the base session and incremental sessions, we develop a dual selective SSM projector to separately exploit the dynamic adaptation. Concretely, it consists of two projection branches and an identity branch. Each projection branch splits feature maps into patches and goes through an SS2D module after position embedding. The first one is used to project the backbone features into the classifier prototypes for base classes and is frozen in the incremental sessions. The other one is introduced since the first incremental session, to learn feature shifts of novel classes into their corresponding classifier prototypes.
To further enhance the dynamic adaptation effect, we propose a class-sensitive selective scan mechanism, which guides the selection pattern to be distinctive between the base classes and incremental classes. First, we use a suppression loss that regularizes the norm of gating for the features from the base-session classes. By doing so, the output feature of the base classes in incremental sessions will be close to the one in the base session, preserving the learned knowledge of the base classes as much as possible to induce stability. Second, we introduce a separation loss that minimizes the similarity of the system matrices (${\mathbf{B}}$, ${\mathbf{C}}$, and $\bm{\Delta}$) between the ones output from the base-class features and the ones output from all the incremental-class features. In this way, the selection in SS2D will be performed in different patterns for the base classes and the incremental classes, leaving sufficient feature space for the new classes in the incremental training process to induce plasticity.
To the best of our knowledge, we are the first to explore the potential of Mamba in the context of FSCIL. The contributions of this paper can be summarized as follows:
  • We successfully exploit the potential of Mamba for FSCIL by proposing a dual selective SSM projector. The dual design maintains the base-class features and adaptively adjusts its projection parameters based on the input context for dynamic adaptation on novel classes, mitigating catastrophic forgetting and enhancing adaptability.
  • We further develop a class-sensitive selective scan mechanism to guide the dynamic adaptation composed of two loss terms. The suppression loss minimizes the disruption to old-class representations caused by incremental training, maintaining stability for base classes. The separation loss makes the selective scan perform in distinct patterns to induce plasticity for novel classes.
  • Our approach is thoroughly evaluated on standard FSCIL benchmarks, including CIFAR-100, CUB-200, and miniImageNet. Our results achieve state-of-the-art performances, demonstrating effectiveness in improving and preserving base-class accuracy and enhancing the adaptability to novel classes.
2 Related Work
2.1 Few-Shot Learning (FSL)
FSL enables models to classify new classes with limited training samples [40, 41, 42, 43], and can be divided into four main strategies: data augmentation, metric-based, model-based, and optimization-based approaches [44]. Data augmentation methods reduce overfitting and enhance generalization by increasing data diversity through transforming existing samples or synthetic data generation using models like VAEs or GANs [45]. Metric-based methods classify objects by computing similarity or distance between samples in a learned embedding space, with common approaches including Siamese Networks [46], Matching Networks [47], and Prototypical Networks [48]. Model-based methods design specific architectures for FSL challenges, such as Memory-Augmented Neural Networks[49] and Meta Networks [50], to store crucial class information and learn generalizable meta-knowledge. Optimization-based methods aim for fast model adaptation with few-shot data. Model-Agnostic Meta-Learning (MAML) [51] is a common optimization algorithm that trains the model’s initial parameters using various datasets to ensure a peak performance when tackling new tasks. Reptile [52] extends MAML by simplifying gradient calculations, enhancing computational efficiency.
While FSL focuses on generalizing to new classes with limited data, it does not prioritize retaining knowledge of base classes. In contrast, FSCIL requires models to continuously learn new classes with limited samples while preserving the knowledge of previously learned classes.
2.2 Class-Incremental Learning (CIL)
CIL enables models to learn new classes over time without forgetting previously acquired knowledge [53, 54, 55, 56, 57, 9, 24, 58]. CIL approaches include non-exemplar-based and exemplar-based methods. Non-exemplar-based methods focus on preserving existing knowledge without relying on stored examples. Methods like EWC [59], MAS [60], and SI [61] measure the importance of parameters and apply regularization to prevent significant changes in critical parameters. Additionally, some methods dynamically change the model architecture, using network expansion and pruning to accommodate new tasks [62, 63, 22]. Exemplar-based methods, e.g., iCaRL [55], combine exemplar replay with knowledge distillation, while BiC [64] forms a validation set from saved exemplars and adjusts outputs via a scaling layer. GEM [65, 66] uses exemplars for gradient projection to prevent knowledge overwriting. Additional approaches include storing embeddings rather than raw images [67], using generative models for replay [68], and implementing task-wise adaptation and output normalization [69, 70, 71, 72].
While CIL typically has sufficient training data in incremental sessions, FSCIL complicates this by limiting training samples, making the prevention of catastrophic forgetting even more difficult.
2.3 Few-shot Class-Incremental Learning (FSCIL)
FSCIL [4, 5] combines the aspects of FSL and CIL, continuously learning new classes with a small number of samples to a model’s knowledge base over time without retraining from scratch. FSCIL faces the key challenges: (1) Catastrophic Forgetting. Learning new classes without accessing the data of old classes can cause the model to forget previously acquired knowledge  [73, 74, 75]. (2) Data Scarcity. Limited data availability hampers the model’s ability to learn robust and generalizable representations and increases the overfitting risk [48, 76]. (3) Stability-Plasticity Dilemma. The model must seek a delicate balance, maintaining the knowledge of old classes while adapting to the new ones [77, 78, 79].
Various methods have been proposed to address the challenges. Data-based strategies mitigate catastrophic forgetting by replaying a subset of real data or generating synthetic data [6, 7, 8]. Optimization-based methods adopt techniques like meta-learning [20, 17, 18], pseudo-feature generation [80, 3], advanced loss functions or regularizers [9, 14, 15, 11, 12, 81, 82, 83, 84], well-designed classifiers [16, 19, 85, 86, 87], and knowledge distillation [1, 13, 10, 88]. For example, C-FSCIL [16] employs a mapping strategy that transforms input images into quasi-orthogonal prototypes to enhance class separability. NC-FSCIL [19] introduces a fixed equiangular tight frame classifier and trains a projection layer to attract the backbone features to these ideal prototypes, thus mitigating inter-class interference. However, most of these methods are based on static adaptation, whose parameters are not aware of the input, increasing the difficulty of striving for a comprehensive ability in both old and new classes. Dynamic adaptation methods dynamically adjust the model structure or interrelationships between prototypes [1, 21, 22, 23, 2, 89]. TOPIC [1] uses a neural gas network to integrate new nodes and edges for new classes. CEC [2] employs graph models to evolve the classifier’s topology as new classes emerge. Dynamic networks such as LEC-Net [21], DER [22], DSN [23] and FeSSS [89] introduce task-specific parameters for new tasks while preserving those related to old tasks.
Different from these approaches that often require expanding the parameter space, our method achieves dynamic adaptation leveraging the Selective SSM mechanism in Mamba that adjusts parameters based on input characteristics without increasing model complexity.
2.4 State Space Models
Transformers [28, 29, 30], despite their proficiency in handling long-range dependencies, are hindered by their quadratic computational cost with respect to input size and large inference-time memory requirements from the key-value cache, particularly for long sequences or high-resolution images. In contrast, State Space Models (SSMs) provide an efficient alternative, capable of processing extensive sequences with significantly reduced computational demands [90, 91, 92, 93, 94, 95, 96, 97, 98, 99]. SSMs involve recurrent updates over a sequence through hidden states, maintaining context in hidden states, and updating outputs by integrating these states with incoming inputs. Structured state space models (S4) [90, 91] exemplify this efficiency with a parameterized linear recurrent neural network. Subsequent research further enhances computational efficiency and expressive capacity of SSMs across various applications [91, 92, 93, 94, 95].
Mamba [27, 100] enhances S4 with a data-dependent selective mechanism to selectively propagate or forget information according to sequential input tokens and proposes hardware-aware algorithms to boost efficiency. Mamba has also been adopted to process non-sequential input such as images and videos in vision tasks  [31, 35, 33, 34, 37, 38, 36, 39]. Among them, VMamba [31] introduces a cross-scan mechanism that processes images in both horizontal and vertical dimensions, allowing feature map elements to integrate information from all directions and providing a global receptive field without increasing linear computational complexity.
Although the Mamba model has been extensively applied in various domains such as text [27, 96, 90, 101], vision [35, 31, 36, 102], audio [103], its potential to mitigate model forgetting has not yet been explored. This study is the first attempt to apply Mamba in FSCIL, effectively harnessing its capabilities of dynamic adaptation and sequence modeling to achieve state-of-the-art performance.
3 Preliminaries
3.1 Few-Shot Class-Incremental Learning
FSCIL trains a model incrementally in multiple sessions $\{{\mathcal{D}}^{(0)},{\mathcal{D}}^{(1)},\dots,{\mathcal{D}}^{(T)}\}$, where ${\mathcal{D}}^{(t)}=\{({\bm{x}}_{i},y_{i})\}_{i=1}^{|{\mathcal{D}}^{(t)}|}$ represents the training set for session $t$. ${\mathcal{D}}^{(0)}$ is the base session, and $T$ is the number of incremental sessions. The base session ${\mathcal{D}}^{(0)}$ usually contains a large label space ${\mathcal{C}}^{(0)}$ and extensive training data for each class $c\in{\mathcal{C}}^{(0)}$. In each incremental session ${\mathcal{D}}^{(t)}$, $t>0$, there are only a few labeled images, $|{\mathcal{D}}^{(t)}|=pq$, where $p$ is the number of classes and $q$ is the number of samples per novel class, known as $p$-way $q$-shot. There is no overlap in the label space between sessions i.e., ${\mathcal{C}}^{(t)}\cap{\mathcal{C}}^{(t^{\prime})}=\emptyset$ for all $t^{\prime}\neq t$. The training sets from previous sessions are not accessible, which requires the model to generalize to new data without forgetting previously learned information. Evaluation in session $t$ involves test data from all classes encountered up to that session, i.e., the label space of $\cup_{i=0}^{t}{\mathcal{C}}^{(i)}$.
3.2 Mamba
State Space Models (SSMs) map the input sequence $x(t)\in\mathbb{R}^{L}$ to the output sequences $y(t)\in\mathbb{R}^{L}$ through a recurrently updated hidden state $h(t)\in\mathbb{R}^{D}$, formulated as:
$h^{\prime}(t)={\mathbf{A}}h(t)+{\mathbf{B}}x(t),\quad y(t)={\mathbf{C}}h(t),$ (1)
where ${\mathbf{A}}\in\mathbb{R}^{D\times D}$ is the system’s evolution matrix, and ${\mathbf{B}}\in\mathbb{R}^{D\times 1}$, ${\mathbf{C}}\in\mathbb{R}^{1\times D}$ are the projection matrix. To adapt this continuous system for practical implementation, the structured state space model (S4) [90] and Mamba [27] employ the discretization transformation by zero-order hold rule, using a timescale parameter $\mathbf{\Delta}\in\mathbb{R}>0$ to transform the continuous parameters to discrete counterparts:
$\overline{{\mathbf{A}}}$ $=\exp{(\mathbf{\Delta}{\mathbf{A}})},$ (2)
$\overline{{\mathbf{B}}}$ $=(\mathbf{\Delta}{\mathbf{A}})^{-1}(\exp{(\mathbf{\Delta}{\mathbf{A}})}-\mathbf{I})\cdot\mathbf{\Delta}{\mathbf{B}}.$
The discretized system is defined as:
$h_{t}=\overline{{\mathbf{A}}}h_{t-1}+\overline{{\mathbf{B}}}x_{t},\quad y_{t}={\mathbf{C}}h_{t}.$ (3)
This system enables efficient global convolution operations that can be processed in parallel:
$\overline{{\mathbf{K}}}$ $=({\mathbf{C}}\overline{{\mathbf{B}}},{\mathbf{C}}\overline{{\mathbf{A}}}\overline{{\mathbf{B}}},\dots,{\mathbf{C}}\overline{{\mathbf{A}}}^{L-1}\overline{{\mathbf{B}}}),\quad\mathbf{y}={\mathbf{X}}*\overline{{\mathbf{K}}},$ (4)
where * denotes convolution operation, and $\overline{{\mathbf{K}}}\in\mathbb{R}^{L}$ is a convolutional kernel that covers the entire input sequence with the length of $L$. However, the current system employs constant operation parameters, e.g., ${{\mathbf{C}}}$ and ${{\mathbf{B}}}$, which remain static for varying input. This mechanism prevents S4 from selecting the correct information from the context or affecting the hidden state in an input-dependent manner.
Selective State Space Models (S6) are the core component of Mamba, enhancing S4 by introducing a selective and dynamic mechanism for the interactions between sequential states [27]. Unlike conventional SSMs that utilize constant operation matrices that are agnostic of input features, S6 models utilize input-dependent parameters that calculate parameters ${\mathbf{B}}$, ${\mathbf{C}}$, and $\mathbf{\Delta}$ directly from the input sequence ${\bm{x}}$ through linear projections, enabling a dynamic and context-aware adaptation to the unique characteristics of the data.
2D Selective Scan (SS2D) enhances Mamba’s ability to effectively capture spatial information for vision tasks [31]. SS2D arranges the feature patches in four different directions (top-left to bottom-right, bottom-right to top-left, top-right to bottom-left, and bottom-left to top-right), creating four separate sequences. Each sequence is processed by the S6 module before merging to reconstruct the 2D feature map. Given the input feature ${\bm{z}}$, the output feature $\overline{{\bm{z}}}$ of SS2D is formulated as:
$\overline{{\bm{z}}}=\text{SS2D}({\bm{z}})=\sum_{i=1}^{K}\text{S6}(\text{scan}({\bm{z}},i)),$ (5)
where $\text{scan}({\bm{z}},i)$ converts feature patches into a sequence according to direction $i$, and $K$ denotes the number of scan paths (default is 4). This approach ensures the comprehensive integration of contextual information from all directions, enhancing the model’s global receptive field.
4 Mamba-FSCIL
Figure 2: Mamba-FSCIL framework. The macro architecture (a) is composed of a backbone network $f$, a dual selective SSM projector consisting of an identity branch $p$ and two selective SSM branches ($g^{\text{base}}$ for the base session and $g^{\text{inc}}$ for incremental sessions) with the identical structure of (b), and a classifier. Our class-sensitive selective scan (c) performs $\mathcal{L}_{\text{supp}}$ and $\mathcal{L}_{\text{sep}}$ in the incremental selective SSM branch $g^{\text{inc}}$ to guide the dynamic adaptation in incremental session’s training.
4.1 Macro Architecture
The architecture of Mamba-FSCIL, illustrated in Fig. 2 (a), is composed of a backbone network $f$, a dual selective SSM projector, and a linear classifier. The backbone network $f$ processes the input ${\bm{x}}$ to generate intermediate features ${\mathbf{F}}=f({\bm{x}})\in\mathbb{R}^{N\times D\times H\times W}$, where $N$, $D$, $H$, and $W$ denote the batch size, number of channels, height, and width, respectively. It is only trainable in the base session to learn a powerful feature extractor from the extensive base session data and is then frozen during incremental sessions, a common practice in FSCIL [2, 19, 104, 83] to prevent catastrophic forgetting caused by fine-tuning on novel classes.
We develop a dual selective SSM projector that maps the frozen intermediate features into the final representation for dynamic adaptation. It consists of three branches: a base session branch $g^{\text{base}}$, an incremental session branch $g^{\text{inc}}$, and an identity branch $p$. The base session branch and incremental session branch have the same structure shown in Fig. 2 (b). The base session branch is responsible for projecting the features of the base classes into a stable and robust representation, which is frozen during the incremental sessions. The incremental session branch dynamically adapts to new classes by learning feature shifts while preserving the previously learned knowledge. The details of this projector will be introduced in Section 4.2.
To further enhance dynamic adaptation, a class-sensitive selective scan mechanism is proposed. This mechanism guides the incremental session branch $g^{\text{inc}}$ in adapting to new classes with minimal interference to the representations of base classes. As shown in Fig. 2 (c), it employs suppression and separation losses to manage the impact of new classes and enforce distinct patterns of feature adaptation between base and novel classes. This mechanism will be detailed in Section 4.3.
The classifier adopts a fixed simplex equiangular tight frame (ETF) following [19, 105, 106, 107, 108, 109]. It provides an ideal feature space arrangement with the maximal equiangular separation among classes [110, 111]. In Section 4.4, we will introduce the training and optimization process across sessions.
4.2 Dual Selective SSM Projector
As illustrated in Figure 2, a backbone network produces the intermediate feature ${\mathbf{F}}$. The dual selective SSM projector is composed of three branches: an identity branch, and two selective SSM branches. The identity branch processes ${\mathbf{F}}$ through ${\bm{\mu}}^{\text{iden}}=p(\text{AvgPool}({\mathbf{F}}))$, where AvgPool is the average pooling function, $p$ is linear projection operation. In the selective SSM branch for the base session ($g^{\text{base}}$), as shown in Fig. 2 (b), ${\mathbf{F}}$ is first reshaped from a 2D feature map into a sequence of flattened feature vectors to be processed by an MLP layer and is added with the position embedding as:
$\hat{{\mathbf{F}}}^{\text{base}}=\rm{MLP}(\text{Reshape}({\mathbf{F}}))+{\mathbf{E}}^{\text{base}},$ (6)
where ${\mathbf{E}}^{\text{base}}\in\mathbb{R}^{L\times D^{\prime}}$ denotes the position embedding, $\hat{{\mathbf{F}}}^{\text{base}}\in\mathbb{R}^{N\times L\times D^{\prime}}$ is the output feature, $D^{\prime}$ denotes the output dimension, and $L=H\times W$ denotes the total number of spatial locations. Then $\hat{{\mathbf{F}}}^{\text{base}}$ goes through a selective SSM module, which splits the features into two information streams via linear projections $f_{x}$ and $f_{z}$, creating ${\mathbf{X}}^{\text{base}}=f_{x}(\hat{{\mathbf{F}}}^{\text{base}})$ and ${\mathbf{Z}}^{\text{base}}=f_{z}(\hat{{\mathbf{F}}}^{\text{base}})$. ${\mathbf{X}}^{\text{base}}$ is intended for scanning, while ${\mathbf{Z}}^{\text{base}}$ is used for gating the output of the scanned results. Specifically, the ${\mathbf{X}}^{\text{base}}$ stream is then enhanced through a group convolution function followed by a $\rm{SiLU}$ activation function:
$\hat{{\mathbf{X}}}^{\text{base}}=\rm{SiLU}(\text{Conv}({\mathbf{X}}^{\text{base}})).$ (7)
The enhanced features $\hat{{\mathbf{X}}}^{\text{base}}$ are then processed by the SS2D module defined in Eq. (5), enhancing discrimination and stability crucial for maintaining high accuracy within base classes. The output of the selective SSM branch ${\bm{\mu}}^{\text{base}}$ is obtained by elementwise multiplication of the output of the SS2D module and ${\mathbf{Z}}^{\text{base}}$ (after applying the $\rm{SiLU}$ activation) with average pooling:
${\bm{\mu}}^{\text{base}}=\text{AvgPool}\left(\text{SS2D}(\hat{{\mathbf{X}}}^{\text{base}})\otimes\text{SiLU}({\mathbf{Z}}^{\text{base}})\right),$ (8)
where $\otimes$ denotes elementwise multiplication. The SS2D system matrices for the input features $\hat{{\mathbf{X}}}^{\text{base}}$ are calculated by linear projections as:
${\mathbf{B}}=f_{B}(\hat{{\mathbf{X}}}^{\text{base}}),\quad{\mathbf{C}}=f_{C}(\hat{{\mathbf{X}}}^{\text{base}}),\quad\bm{\Delta}=f_{\Delta}(\hat{{\mathbf{X}}}^{\text{base}}),$ (9)
where ${\mathbf{B}}\in\mathbb{R}^{N\times K\times D_{B}\times L}$, ${\mathbf{C}}\in\mathbb{R}^{N\times K\times D_{C}\times L}$, $\bm{\Delta}\in\mathbb{R}^{N\times K\times D_{\Delta}\times L}$ are data-dependent system matrices, with $K$ representing the number of scan directions in Eq. (5). Here, $D_{*}$ represents the dimensions of each parameter. Then, the final representation of the base session is obtained by combining the outputs of the identity branch and the selective SSM branch as:
${\bm{\mu}}^{(0)}={\bm{\mu}}^{\text{base}}+{\bm{\mu}}^{\text{iden}}.$ (10)
During the base session, the backbone network and the selective SSM branch $g^{\text{base}}$ use the abundant available data to learn robust feature representations, thus laying a powerful feature extractor for later incremental training. Due to the huge gap in training data between the base session and incremental sessions, a single selective SSM is difficult to achieve a universal dynamic adaptation for all the base and incremental classes. To this end, our dual selective SSM projector adopts another selective SSM branch ($g^{\text{inc}}$) with an identical structure for incremental sessions.
Concretely, in each incremental session $1\leq t\leq T$, the backbone network $f$, the identity branch $p$, and the selective SSM branch for the base session $g^{\text{base}}$ are frozen, and only the newly introduced $g^{\text{inc}}$ is learnable to adapt the model to novel classes. Following the widely adopted practice in FSCIL [10, 11, 12, 16, 19, 82], we employ a memory module ${\mathcal{M}}^{(t)}$, which stores the mean intermediate feature ${\mathbf{F}}_{c}$ of each old class $c$ to alleviate forgetting when encountering novel classes:
${\mathcal{M}}^{(t)}=\{{\mathbf{F}}_{c}\mid c\in\cup_{j=0}^{t-1}{\mathcal{C}}^{(j)}\},$ (11)
$\quad{\mathbf{F}}_{c}=\mathrm{Avg}_{i}\{f({\bm{x}}_{i})\mid y_{i}=c\}.$
The intermediate features of the new class data, along with the stored mean features of the old classes, are processed through $p$ and $g^{\text{base}}$ to produce ${\bm{\mu}}^{\text{iden}}$ and ${\bm{\mu}}^{\text{base}}$ respectively, using the same processes as in the base session. Additionally, these features are processed through $g^{\text{inc}}$ to generate the output:
${\bm{\mu}}^{\text{inc}}=\text{AvgPool}\left(\text{SS2D}(\hat{{\mathbf{X}}}^{\text{inc}})\otimes\text{SiLU}({\mathbf{Z}}^{\text{inc}})\right),$ (12)
where $\hat{{\mathbf{X}}}^{\text{inc}}$ and ${\mathbf{Z}}^{\text{inc}}$ are processed using the same architecture as the base branch $g^{\text{base}}$, but with different parameters. The resulting final representation for incremental sessions is aggregated as:
${\bm{\mu}}^{\text{(t)}}={\bm{\mu}}^{\text{base}}+{\bm{\mu}}^{\text{inc}}+{\bm{\mu}}^{\text{iden}},\quad t\geq 1.$ (13)
In this way, the branch $g^{\text{inc}}$ learns the shifts of feature distribution prompted by the new class data upon the robust representation established during the base session to dynamically accommodate the model to new classes. It enhances the adaptation ability of the model without altering the previously learned knowledge badly under the guidance of our class-sensitive selective scan.
4.3 Class-sensitive Selective Scan
In this section, we develop a class-sensitive selective scan mechanism, as illustrated in Fig. 2 (c) that guides the dynamic adaptation of the incremental selective SSM branch $g^{\text{inc}}$ by two loss terms: suppression loss (${\mathcal{L}}_{\text{supp}}$) and separation loss (${\mathcal{L}}_{\text{sep}}$).
Suppression Loss (${\mathcal{L}}_{\text{supp}}$). This loss focuses on adjusting the norm of the feature vector $\mathbf{Z}^{\text{inc}}$, depicted in the top part of Fig. 2 (c). The diagram shows the distribution of features for the base (blue dots) and novel (orange dots) classes, with dashed lines representing the intention of modifying feature norms. In Eq. (12), the feature vector ${\mathbf{Z}}^{\text{inc}}$ plays a pivotal role in the elementwise multiplication with outputs from the SS2D module. By adjusting the norm of ${\mathbf{Z}}^{\text{inc}}$, we intend to manipulate the magnitude of the resulting feature vector ${\bm{\mu}}^{\text{inc}}$, which serves as a class-sensitive gating mechanism that controls the impact of new features with respect to base classes and incremental classes differently during incremental session training:
  • For base classes, the norm of ${\mathbf{Z}}^{\text{inc}}$ of base classes is suppressed (red dashed lines) to decrease ${\bm{\mu}}^{\text{inc}}$’s magnitude, thus minimizing feature shifts caused by the new branch $g^{\text{inc}}$ to maintain the stability of the features of the base classes.
  • For novel classes, conversely, the norm of ${\mathbf{Z}}^{\text{inc}}$ is increased (green dashed lines) to enhance ${\bm{\mu}}^{\text{inc}}$’s magnitude, facilitating the learning of feature shifts that help the model adapt to new classes.
Therefore, the suppression loss can be formulated as:
${\mathcal{L}}^{\text{base}}_{\text{supp}}$ $=\sum_{({\mathbf{F}}_{c},y_{c})\in{\mathcal{M}}^{(0)}}\|{\mathbf{Z}}^{\text{inc}}_{c}\|^{2},$ (14)
${\mathcal{L}}^{\text{novel}}_{\text{supp}}$ $=-\sum_{({\bm{x}}_{i},y_{i})\in{\mathcal{D}}^{(t)}}\|{\mathbf{Z}}^{\text{inc}}_{i}\|^{2}-\sum_{({\mathbf{F}}_{c},y_{c})\in{{\mathcal{M}}}^{(t)}\backslash{{\mathcal{M}}}^{(0)}}\|{\mathbf{Z}}^{\text{inc}}_{c}\|^{2},$
where ${{\mathcal{M}}}^{(t)}\backslash{{\mathcal{M}}}^{(0)}=\{{\mathbf{F}}_{c}\mid c\in\cup_{j=1}^{t-1}{\mathcal{C}}^{(j)}\}$, and we have:
${\mathcal{L}}_{\text{supp}}={\mathcal{L}}^{\text{base}}_{\text{supp}}+{\mathcal{L}}^{\text{novel}}_{\text{supp}}.$ (15)
By differentially managing the feature magnitudes for base and novel classes through suppression loss, the interference with base-class features brought by the newly added branch $g_{\text{inc}}$ is diminished, ensuring that new class integration does not destabilize the model’s existing knowledge. In the ideal case, it is expected that the output of $g^{\text{inc}}$, i.e., ${\bm{\mu}}^{\text{inc}}$, could be zero such that the inference of base-class data after incremental session’s training (Eq. (13)) will be equivalent to the one after the base session’s training (Eq. (10)) to induce stability.
Separation Loss (${\mathcal{L}}_{\text{sep}}$). The separation term is designed to enhance class sensitivity in the input-dependent parameters (${\mathbf{B}},{\mathbf{C}},\bm{\Delta}$) within the SS2D module, ensuring that base and novel classes are effectively distinguished. As illustrated in the lower part of Fig. 2 (c), the separation mechanism ”pushes” the averaged input-dependent parameters associated with novel classes (${\mathbf{B}}_{n},{\mathbf{C}}_{n},\bm{\Delta}_{n}$) away from those of the base classes (${\mathbf{B}}_{b},{\mathbf{C}}_{b},\bm{\Delta}_{b}$), thereby maximizing the distinctiveness between these classes.
Specifically, for both base and novel classes, we calculate their average input-dependent parameters:
$\bar{{\mathbf{P}}}_{\text{b}}$ $=\frac{1}{N_{\text{base}}\cdot K\cdot L}\sum_{({\mathbf{F}}_{i},y_{i})\in{\mathcal{M}}^{(0)}}\sum_{j=1}^{K}\sum_{k=1}^{L}{\mathbf{P}}_{i,j,:,k},$ (16)
$\bar{{\mathbf{P}}}_{\text{n}}$ $=\frac{1}{N_{\text{novel}}\cdot K\cdot L}\sum_{{\mathcal{D}}^{(t)}\cup({{\mathcal{M}}}^{(t)}\backslash{{\mathcal{M}}}^{(0)})}\sum_{j=1}^{K}\sum_{k=1}^{L}{\mathbf{P}}_{i,j,:,k},$
where $\bar{{\mathbf{P}}}_{\text{b}}$ and $\bar{{\mathbf{P}}}_{\text{n}}$ represent the aggregated parameters for the base and novel classes, respectively. The symbol ${\mathbf{P}}$ represents one of ${\mathbf{B}}$, ${\mathbf{C}}$, or $\bm{\Delta}$. The terms $N_{\text{base}}$ and $N_{\text{novel}}$ denote the number of samples in base and novel classes in a mini-batth. To maximize separability, we then compute and minimize the cosine similarity of these averaged parameters between base and novel classes ($\bar{{\mathbf{P}}}_{\text{b}}$ and $\bar{{\mathbf{P}}}_{\text{n}}$), which can be formulated as:
${\mathcal{L}}_{\text{sep}}=\frac{\bar{{\mathbf{B}}}_{\text{b}}\cdot\bar{{\mathbf{B}}}_{\text{n}}}{\|\bar{{\mathbf{B}}}_{\text{b}}\|\|\bar{{\mathbf{B}}}_{\text{n}}\|}+\frac{\bar{{\mathbf{C}}}_{\text{b}}\cdot\bar{{\mathbf{C}}}_{\text{n}}}{\|\bar{{\mathbf{C}}}_{\text{b}}\|\|\bar{{\mathbf{C}}}_{\text{n}}\|}+\frac{\bar{\bm{\Delta}}_{\text{b}}\cdot\bar{\bm{\Delta}}_{\text{n}}}{\|\bar{\bm{\Delta}}_{\text{b}}\|\|\bar{\bm{\Delta}}_{\text{n}}\|}.$ (17)
By minimizing ${\mathcal{L}}_{\text{sep}}$, the selective scan can be performed in distinct patterns between base and novel classes to induce plasticity during incremental learning.
4.4 Optimization
After we get the final representation, i.e., Eq. (10) for the base session and Eq. (13) for incremental sessions, we calculate the classification loss with the ETF classifier. Following [19, 107], we use the dot regression (DR) loss [110] as the training objective, as it is shown to outperform cross entropy loss when using a fixed ETF classifier in imbalanced training, which FSCIL naturally satisfies. The DR loss can be written as:
${\mathcal{L}}_{\text{DR}}\left(\hat{{\bm{\mu}}}_{i},\hat{{\mathbf{W}}}_{\rm ETF}\right)=\frac{1}{2}\left(\hat{\mathbf{w}}^{T}_{y_{i}}\hat{{\bm{\mu}}}_{i}-1\right)^{2},$ (18)
where $\hat{{\bm{\mu}}}_{i}=\frac{{\bm{\mu}}_{i}}{\|{\bm{\mu}}_{i}\|}$ is the $l_{2}$ normalized feature of the final representation, $y_{i}$ is the label of input $x_{i}$, and $\hat{\mathbf{w}}_{y_{i}}$ is the classifier prototype in $\hat{{\mathbf{W}}}_{\rm ETF}$ for label $y_{i}$.
Base Session Training. During base session training ($t=0$), the backbone network $f$, the identity branch $p$, and the selective SSM branch for the base session $g^{\text{base}}$ are jointly optimized to establish a robust feature extractor. The objective during this phase is to minimize the classification loss on the extensive base session dataset:
$\min_{f,p,g^{\text{base}}}\frac{1}{|{\mathcal{D}}^{(0)}|}\sum_{({\bm{x}}_{i},y_{i})\in{\mathcal{D}}^{(0)}}{\mathcal{L}}_{\text{DR}}\left(\hat{{\bm{\mu}}}_{i}^{(0)},\hat{{\mathbf{W}}}_{\rm ETF}\right).$ (19)
Incremental Sessions Training. In subsequent sessions ($1\leq t\leq T$), as new classes are introduced, the optimization of the model focuses on seamlessly integrating these classes while minimizing disruption of previously learned knowledge. The training objective during these sessions is to adapt the model with new-class data and maintain the existing ability of old classes with the stored intermediate features in ${\mathcal{M}}^{(t)}$. Thus, the classification loss can be formulated as:
${\mathcal{L}}_{\text{cls}}^{(t)}=\frac{1}{|{\mathcal{D}}^{(t)}|+|{\mathcal{M}}^{(t)}|}\left({\mathcal{L}}^{(t)}_{\text{cls-new}}+{\mathcal{L}}^{(t)}_{\text{cls-old}}\right),$ (20)
where
${\mathcal{L}}^{(t)}_{\text{cls-new}}$ $=\sum_{({\bm{x}}_{i},y_{i})\in{\mathcal{D}}^{(t)}}{\mathcal{L}}_{\text{DR}}\left(\hat{{\bm{\mu}}}_{i}^{(t)},\hat{{\mathbf{W}}}_{\rm ETF}\right),$ (21)
${\mathcal{L}}^{(t)}_{\text{cls-old}}$ $=\sum_{({\mathbf{F}}_{c},y_{c})\in{\mathcal{M}}^{(t)}}{\mathcal{L}}_{\text{DR}}\left(\hat{{\bm{\mu}}}_{c}^{(t)},\hat{{\mathbf{W}}}_{\rm ETF}\right).$
Together with the suppression loss and the separation loss in Section 4.3, the total objective in session $t\geq 1$ is:
$\min_{g^{\text{inc}}}\left({\mathcal{L}}_{\text{cls}}^{(t)}+\alpha{\mathcal{L}}_{\text{supp}}^{(t)}+\beta{\mathcal{L}}_{\text{sep}}^{(t)}\right),$ (22)
where $\alpha$ and $\beta$ are hyper-parameters to balance the loss scales.
5 Experiments
Our experiments evaluate Mamba-FSCIL on the three widely used benchmarks, including miniImageNet [112], CIFAR-100 [113], and CUB-200 [114], comparing it with static and dynamic methods. We conducted ablation studies to analyze the contributions of key components, including the dual selective SSM projector and the class-sensitive selective scan mechanism, as well as the effects of suppression and separation losses on feature alignment and class separability. We also examined the impact of freezing versus training the identity and selective SSM branches, the sensitivity to hyper-parameters, and the influence of varying training samples (shots) and scan paths ($K$). Additionally, we compared different projection architectures and evaluated Mamba-FSCIL against Dynamic Support Networks (DSN) [23] to demonstrate its efficiency in managing model complexity while achieving superior performance. Finally, we visualized activation maps to further validate the effectiveness of our approach and provided an analysis of feature embeddings.
TABLE I: FSCIL performance across sessions on miniImageNet compared with other methods. ”Average Acc.” is the average accuracy across all sessions. ”Final Improv.” denotes the accuracy improvement of our method in the last session compared to previous methods.
Methods Venue Accuracy in each session (%) $\uparrow$ Average Final
0 1 2 3 4 5 6 7 8 Acc. Improv.
Static adaptation methods
IDLVQ [11] ICLR 2021 64.77 59.87 55.93 52.62 49.88 47.55 44.83 43.14 41.84 51.16 +17.52
Self-promoted [86] CVPR 2021 61.45 63.80 59.53 55.53 52.50 49.60 46.69 43.79 41.92 52.76 +17.44
SFbFSCIL [80] ICCV 2021 61.40 59.80 54.20 51.69 49.45 48.00 45.20 43.80 42.10 50.63 +17.26
Data-free Replay [6] ECCV 2022 71.84 67.12 63.21 59.77 57.01 53.95 51.55 49.52 48.21 58.02 +11.15
LIMIT [18] TPAMI 2022 72.32 68.47 64.30 60.78 57.95 55.07 52.70 50.72 49.19 59.06 +10.17
MetaFSCIL [17] CVPR 2022 72.04 67.94 63.77 60.29 57.58 55.16 52.90 50.79 49.19 58.85 +10.17
FACT [3] CVPR 2022 72.56 69.63 66.38 62.77 60.6 57.33 54.34 52.16 50.49 60.70 +8.87
C-FSCIL [16] CVPR 2022 76.40 71.14 66.46 63.29 60.42 57.46 54.78 53.11 51.41 61.61 +7.95
Regularizer [12] ICLR 2022 80.37 74.68 69.39 65.51 62.38 59.03 56.36 53.95 51.73 63.71 +7.63
TEEN [87] NeurIPS 2023 73.53 70.55 66.37 63.23 60.53 57.95 55.24 53.44 52.08 61.44 +7.28
CABD [88] CVPR 2023 74.65 70.43 66.29 62.77 60.75 57.24 54.79 53.65 52.22 61.42 +7.14
ALICE [8] ECCV 2022 80.60 70.60 67.40 64.50 62.50 60.00 57.80 56.80 55.70 63.99 +3.66
NC-FSCIL [19] ICLR 2023 84.02 76.80 72.00 67.83 66.35 64.04 61.46 59.54 58.31 67.82 +1.05
Dynamic adaptation methods
TOPIC [1] CVPR 2020 61.31 50.09 45.17 41.16 37.48 35.52 32.19 29.46 24.42 39.64 +34.94
LEC-Net [21] arXiv 2022 61.31 35.37 36.66 38.59 33.90 35.89 36.12 32.97 30.55 37.92 +29.69
DSN [23] TPAMI 2022 68.95 63.46 59.78 55.64 52.85 51.23 48.9 46.78 45.89 54.83 +13.47
CEC [2] CVPR 2021 72.00 66.83 62.97 59.43 56.70 53.73 51.19 49.24 47.63 57.75 +11.73
FeSSSS [89] CVPR 2022 81.50 77.04 72.92 69.56 67.27 64.34 62.07 60.55 58.87 68.23 +0.49
Mamba-FSCIL (ours) 84.93 80.02 74.61 71.33 69.15 65.62 62.38 60.93 59.36 69.81
5.1 Implementation Details
Datasets. We employed three benchmark datasets. miniImageNet is a variant of ImageNet with 100 classes, each having 500 training and 100 testing images of $84\times 84$ pixels. CIFAR-100 has the same number of classes and images, and the image size is $32\times 32$. CUB-200 is a fine-grained classification dataset consisting of 11,788 images in 200 classes, with an image resolution of $224\times 224$. We followed the standard experimental settings in FSCIL [1, 2, 19]. For miniImageNet and CIFAR-100, the base session includes 60 classes, followed by 8 incremental sessions, each with a 5-way 5-shot setting (5 classes and 5 images per class). For CUB-200, the base session comprises 100 classes, with 10 incremental sessions, each following a 10-way 5-shot setting.
Training Details. We employed the standard data preprocessing and augmentation techniques, including random resizing, flipping, and color jittering [1, 2, 8, 19]. We conducted training with a batch size of 512 for base sessions and 64 for incremental sessions, which includes new session data along with intermediate features stored in memory.
For miniImageNet, we utilized ResNet12 [115] as the backbone and train for 500 epochs in the base session, followed by 100-170 iterations in each incremental session. The initial learning rate is 0.25 for the base session and 0.01 for the incremental sessions.
For CIFAR-100, we utilized ResNet12 [115] as the backbone and trained the base session for 200 epochs, followed by 200 iterations per incremental session. The initial learning rate is 0.25 for both the base and incremental sessions.
For CUB-200, we employed two different architectures. In Table III, ResNet18 [115] serves as the backbone, where we trained the model for 80 epochs in the base session and then proceeded with 200-290 iterations for each incremental session. In Table IV, we utilized the Swin Transformer Tiny [30] with an input resolution of $384\times 384$, we trained the model for 80 epochs during the base session followed by 600 iterations per incremental session. Initial learning rates are set at 0.02 for the backbone and 0.2 for the projector during the base session, and are reduced to 0.05 for the incremental sessions.
We apply a cosine annealing strategy for the learning rate and use SGD with a momentum of 0.009 and a weight decay of 0.0005 as the optimizer. The hyper-parameters $\alpha$ and $\beta$ are set between $50$ and $200$, $0.05$ and $0.5$, respectively. The dimensions of the input-dependent parameters in Eq. (9) are set as $D_{B}=D_{C}=D_{\Delta}=256$. The dimensions $D^{\prime}$ in Eq. (6) are set to $512$, $512$, $1024$ for CIFAR-100, CUB-200, and miniImageNet datasets, respectively. All experiments are conducted using 8 NVIDIA A100 GPUs.
Zero Initialization. In our incremental learning process, we fixed the identity branch and selective SSM branch from the base session $g^{\text{base}}$ and introduced a new selective SSM branch $g^{\text{inc}}$ for the incremental sessions. To prevent disruption of pre-existing feature representations, we applied zero initialization to $g^{\text{inc}}$ by setting the projection parameters of $f_{z}$ for the gating features ${\mathbf{Z}}^{\text{inc}}$ to zero, ensuring that its initial outputs (${\bm{\mu}}^{\text{inc}}$) are also zero. This approach ensures that $g^{\text{inc}}$ begins without affecting the model’s learned features from the base session. As training progresses, $g^{\text{inc}}$ is gradually optimized, allowing the model to adapt to new classes by learning feature shifts while preserving the knowledge acquired during the base session.
5.2 Comparison with the State-of-the-art Methods
TABLE II: FSCIL performance across sessions on CIFAR-100 compared with other methods. ”Average Acc.” is the average accuracy across all sessions. ”Final Improv.” denotes the accuracy improvement of our method in the last session compared to previous methods.
Methods Venue Accuracy in each session (%) $\uparrow$ Average Final
0 1 2 3 4 5 6 7 8 Acc. Improv.
Static adaptation methods
Self-promoted [86] CVPR 2021 64.10 65.86 61.36 57.45 53.69 50.75 48.58 45.66 43.25 54.52 +14.26
Data-free Replay [6] ECCV 2022 74.40 70.20 66.54 62.51 59.71 56.58 54.52 52.39 50.14 60.78 +7.37
MetaFSCIL [17] CVPR 2022 74.50 70.10 66.84 62.77 59.48 56.52 54.36 52.56 49.97 60.79 +7.54
C-FSCIL [16] CVPR 2022 77.47 72.40 67.47 63.25 59.84 56.95 54.42 52.47 50.47 61.64 +7.04
LIMIT [18] TPAMI 2022 73.81 72.09 67.87 63.89 60.70 57.77 55.67 53.52 51.23 61.84 +6.28
FACT [3] CVPR 2022 74.60 72.09 67.56 63.52 61.38 58.36 56.28 54.24 52.10 62.24 +5.41
TEEN [87] NeurIPS 2023 74.92 72.65 68.74 65.01 62.01 59.29 57.90 54.76 52.64 63.10 +4.87
ALICE [8] ECCV 2022 79.00 70.50 67.10 63.40 61.20 59.20 58.10 56.30 54.10 63.21 +3.41
CABD [88] CVPR 2023 79.45 75.38 71.84 67.95 64.96 61.95 60.16 57.67 55.88 66.14 +1.63
NC-FSCIL [19] ICLR 2023 82.52 76.82 73.34 69.68 66.19 62.85 60.96 59.02 56.11 67.50 +1.40
Dynamic adaptation methods
TOPIC [1] CVPR 2020 64.10 55.88 47.07 45.16 40.11 36.38 33.96 31.55 29.37 42.62 +28.14
LEC-Net [21] arXiv 2022 64.10 53.23 44.19 41.87 38.54 39.54 37.34 34.73 34.73 43.14 +22.78
CEC [2] CVPR 2021 73.07 68.88 65.26 61.19 58.09 55.57 53.22 51.34 49.14 59.53 +8.37
DSN [23] TPAMI 2022 73.00 68.83 64.82 62.24 59.16 56.96 54.04 51.57 49.35 60.00 +8.16
FeSSSS [89] CVPR 2022 75.35 70.81 66.70 62.73 59.62 56.45 54.33 52.10 50.23 60.92 +7.28
Mamba-FSCIL (ours) 82.80 77.85 73.69 69.67 66.89 63.66 61.48 59.74 57.51 68.14
TABLE III: FSCIL performance across sessions on CUB-200 compared with other methods. ”Average Acc.” is the average accuracy across all sessions. ”Final Improv.” shows the last session’s accuracy improvement of our method over previous methods.
Methods Venue Accuracy in each session (%) $\uparrow$ Average Final
0 1 2 3 4 5 6 7 8 9 10 Acc. Improv.
Static adaptation methods
Self-promoted [86] CVPR 2021 68.68 61.85 57.43 52.68 50.19 46.88 44.65 43.07 40.17 39.63 37.33 49.32 +24.32
SFbFSCIL [80] ICCV 2021 68.78 59.37 59.32 54.96 52.58 49.81 48.09 46.32 44.33 43.43 43.23 51.84 +18.42
Data-free replay [6] ECCV 2022 75.90 72.14 68.64 63.76 62.58 59.11 57.82 55.89 54.92 53.58 52.39 61.52 +9.26
MetaFSCIL [17] CVPR 2022 75.9 72.41 68.78 64.78 62.96 59.99 58.3 56.85 54.78 53.82 52.64 61.93 +9.01
MgSvF [116] TPAMI 2021 72.29 70.53 67.00 64.92 62.67 61.89 59.63 59.15 57.73 55.92 54.33 62.37 +7.32
FACT [3] CVPR 2022 75.90 73.23 70.84 66.13 65.56 62.15 61.74 59.83 58.41 57.89 56.94 64.42 +4.71
IDLVQ [11] ICLR 2021 77.37 74.72 70.28 67.13 65.34 63.52 62.10 61.54 59.04 58.68 57.81 65.23 +3.84
LIMIT [18] TPAMI 2022 76.32 74.18 72.68 69.19 68.79 65.64 63.57 62.69 61.47 60.44 58.45 66.67 +3.20
TEEN [87] NeurIPS 2023 77.26 76.13 72.81 68.16 67.77 64.40 63.25 62.29 61.19 60.32 59.31 66.63 +2.34
NC-FSCIL [19] ICLR 2023 80.45 75.98 72.30 70.28 68.17 65.16 64.43 63.25 60.66 60.01 59.44 67.28 +2.21
ALICE [8] ECCV 2022 77.40 72.70 70.60 67.20 65.90 63.40 62.90 61.90 60.50 60.60 60.10 65.75 +1.55
Dynamic adaptation methods
TOPIC [1] CVPR 2020 68.68 62.49 54.81 49.99 45.25 41.40 38.35 35.36 32.22 28.31 26.28 43.92 +35.37
LEC-Net [21] arXiv 2022 70.86 58.15 54.83 49.34 45.85 40.55 39.70 34.59 36.58 33.56 31.96 45.08
CEC [2] CVPR 2021 75.85 71.94 68.50 63.50 62.43 58.27 57.73 55.81 54.83 53.52 52.28 61.33 +9.37
FeSSSS [89] CVPR 2022 79.60 73.46 70.32 66.38 63.97 59.63 58.19 57.56 55.01 54.31 52.98 62.85 +8.67
DSN [23] TPAMI 2022 76.06 72.18 69.57 66.68 64.42 62.12 60.16 58.94 56.99 55.10 54.21 63.31 +7.44
Mamba-FSCIL (ours) 80.90 76.26 72.97 70.14 67.83 65.74 65.43 64.12 62.31 62.12 61.65 68.13
The comparison with the state-of-the-art methods on miniImageNet, CIFAR-100, and CUB-200 is shown in Table I, Table II, and Table III. The results demonstrate that Mamba-FSCIL consistently outperforms almost all existing static and dynamic adaptation methods by over 1% in terms of last-session accuracy, where all classes are taken into account. Specifically, on miniImageNet, Mamba-FSCIL achieves the highest average accuracy of 69.81% and increases the last-session accuracy to 59.36%, which surpasses the challenging baseline NC-FSCIL [19] by 1.99% and 1.05%, respectively. In CIFAR-100, Mamba-FSCIL maintains the best performance in most sessions, with the accuracy of the last session outperforming NC-FSCIL by 1.4%. Notably, Mamba-FSCIL significantly surpasses dynamic methods such as FeSSSS [89] with an average accuracy improvement of 7.22% over FeSSSS. In the more challenging fine-grained CUB-200 dataset, Mamba-FSCIL leads with an average accuracy of 68.13%, outperforming all static and dynamic methods. Compared to DSN, Mamba-FSCIL shows the remarkable ability to learn complex classification scenarios incrementally, improving the accuracy in the last session by 7.44%, also indicating a lower performance drop over sessions.
To further test Mamba-FSCIL on larger vision models, we evaluated it on the CUB-200 dataset using a Swin Transformer-Tiny [30] backbone pre-trained on ImageNet-1K [117]. We compared it against the state-of-the-art methods using the same backbone, including CLOM [104], NC-FSCIL [19], and Comp-FSCIL [83]. As shown in Table IV, Mamba-FSCIL achieves the highest average session accuracy of 78.55% and the lowest performance drop (PD), reducing PD to 14.00% compared to Comp-FSCIL’s 14.87%. Notably, while Comp-FSCIL leverages human-inspired compositional learning for adaptive concept integration, Mamba-FSCIL, without relying on this extra step, still outperforms Comp-FSCIL, demonstrating the strong potential of our approach.
TABLE IV: FSCIL performance across sessions on CUB-200 using Swin Transformer-Tiny backbone. ”Average Acc.” represents the average accuracy across all sessions. ”PD” indicates the performance drop, calculated as the difference in accuracy between the first and last session. Results for CLOM [104] and Comp-FSCIL [83] are taken from Comp-FSCIL [83], while NC-FSCIL [19] results are from our implementation using the same training settings as Mamba-FSCIL for a fair comparison.
Methods Venue Accuracy in each session (%) $\uparrow$ Average PD
0 1 2 3 4 5 6 7 8 9 10 Acc.
CLOM [104] NeurIPS 2022 86.28 82.85 80.61 77.79 76.34 74.64 73.62 72.82 71.24 71.33 70.50 76.18 15.78
NC-FSCIL [19] ICLR 2023 87.53 84.25 81.72 79.10 77.21 75.52 74.51 74.42 72.26 72.86 72.49 77.44 15.04
Comp-FSCIL [83] ICML 2024 87.67 84.73 83.03 80.04 77.73 75.52 74.32 74.55 73.35 73.15 72.80 77.90 14.87
Mamba-FSCIL (ours) 88.13 85.14 83.41 80.77 77.23 75.73 75.70 75.32 74.18 74.26 74.13 78.55 14.00
5.3 Ablation Studies
5.3.1 Impact of Individual Component
TABLE V: Ablation study on the impact of dual selective projector and class-sensitive selective scan mechanism on FSCIL performance across different datasets. “Base” indicates the accuracy in the first session; “Final” refers to the last session; “PD” represents the performance drop between the first and last sessions.
Methods Adaptation miniImageNet CIFAR-100 CUB-200
[…]
Baseline Static 84.02 58.31 25.71 82.52 56.11 26.41 80.45 59.44 21.01
+ Single Selective SSM Projector Dynamic 84.93 57.89 27.04 82.80 56.07 26.73 80.90 60.70 20.20
+ Dual Selective SSM Projector Dynamic 84.93 58.92 26.01 82.80 56.71 26.09 80.90 61.06 19.84
+ Dual Selective SSM Projector + Class-sensitive Selective Scan Dynamic 84.93 59.36 25.57 82.80 57.51 25.29 80.90 61.65 19.25
We performed ablation studies across the three datasets to assess the impact of our dual selective projector and class-sensitive selective scan mechanism. Utilizing a consistent setup that includes the same backbone, identity branch, memory module, classifier, and classification loss, we explored several configurations: (1) Baseline employs NC-FSCIL with a static MLP-based projection layer, serving as a competitive static adaptation method. (2) + Single Selective SSM Projector integrates a single dynamic SSM projector with one projection branch, updated across all sessions. (3) + Dual Selective SSM Projector uses our dual selective SSM projector with two branches, as introduced in Section 4.2. (4) + Dual Selective SSM Projector + Class-sensitive Selective Scan further integrates our class-sensitive selective scan.
The results in Table V show that the introduction of a single selective SSM projector improves base session accuracy across the three datasets (e.g., from 84.02% to 84.93% on miniImageNet and from 82.52% to 82.80% on CIFAR-100), demonstrating the Mamba-based projector helps to improve the performance in the base session, providing a more powerful feature extractor for the subsequent sessions. However, despite these gains, the final performances on miniImageNet and CIFAR-100 both fall below the baseline, implying that a naive replacement of the MLP projector with the selective SSM projector cannot effectively mitigate catastrophic forgetting, increasing the performance drop as a result. In contrast, when our proposed dual selective SSM projector is adopted, the final performance is also effectively improved over the baseline, reducing the performance drop from 27.04% to 26.01% on miniImageNet and from 26.73% to 26.09% on CIFAR-100. This shows that our dual design is capable of adaptively learning feature shifts of novel classes into their corresponding classifier prototypes with fewer disruptions to the learned knowledge of the base classes. Further improvements are achieved with the integration of our class-sensitive scan mechanism into the dual projector, increasing the last-session accuracy from 56.71% to 57.51% on CIFAR-100 and from 61.06% to 61.65% on CUB-200. These findings validate the critical role of each component in Mamba-FSCIL.
5.3.2 Impact of Losses in Class-sensitive Selective Scan
Figure 3: Effect of suppression and separation losses in Class-sensitive Selective Scan mechanism on feature-classifier alignment and base-novel class separability in the miniImageNet test set: (a) Average cosine similarity between base-class features and their corresponding classifiers; (b) Average cosine similarity between base and novel classes.
We investigated the impacts of the suppression and separation losses introduced in Section 4.3 by analyzing the features extracted from the test set of the miniImageNet. First, we measured the average cosine similarity between the final representations and the classifier prototypes across all base classes in Fig. 3, defined as:
$\frac{1}{\sum_{k=1}^{K}N_{k}}\sum_{k=1}^{K}\sum_{i=1}^{N_{k}}\cos\langle{\bm{\mu}}_{k,i},\mathbf{w}_{k}\rangle,$ (23)
where ${\bm{\mu}}_{k,i}$ represents the feature vector of sample $i$ in class $k$, $N_{k}$ is the number of samples in class $k$, $\mathbf{w}_{k}$ is the classifier prototype for class $k$, and $K$ is the number of base classes. To assess the separability between base and novel classes, we calculated the average cosine similarity for any possible pairing of base and novel classes feature vectors as:
$\frac{1}{N_{\text{base}}N_{\text{novel}}}\sum_{i=1}^{N_{\text{base}}}\sum_{j=1}^{N_{\text{novel}}}\cos\langle{\bm{\mu}}_{\text{base},i},{\bm{\mu}}_{\text{novel},j}\rangle,$ (24)
where ${\bm{\mu}}_{\text{base},i}$ and ${\bm{\mu}}_{\text{novel},j}$ are feature vectors of the $i$-th base class sample and $j$-th novel class sample, respectively. The terms $N_{\text{base}}$ and $N_{\text{novel}}$ denote the total number of samples in base and novel classes in the test set of miniImageNet.
Fig. 3 (a) shows that the incorporation of suppression loss ${\mathcal{L}}_{\text{supp}}$ helps to improve base-class alignment between features and classifier prototypes in late sessions, facilitating the maintenance of base-class performance. When the separation loss ${\mathcal{L}}_{\text{sep}}$ is further added, as shown in Fig. 3 (b), the similarities of features between the base-class and novel-class are significantly reduced, which leaves sufficient feature space for incremental training sessions and boosts the novel-class accuracy.
5.3.3 Effect of Freezing vs. Training Projection Branches
Figure 4: Impact of freezing vs. training the identity branch $p$ and the selective SSM branch $g^{\text{base}}$ during incremental sessions on miniImageNet: (a) Accuracy for all classes, (b) Accuracy for base classes.
We assessed the impact of freezing the identity branch $p$ and the selective SSM branch $g^{\text{base}}$ from the base session during incremental learning. Models are trained on the miniImageNet dataset, utilizing a ResNet-12 backbone. Two configurations are compared: one with $p$ and $g^{\text{base}}$ frozen (as described in Section 4.2) and the other one with them trainable throughout the incremental sessions. The results in Fig. 4 indicate that the frozen configuration achieves higher accuracy across all classes (Fig. 4 (a)) and specifically for base classes (Fig. 4 (b)) in most sessions. This result implies that freezing $p$ and $g^{\text{base}}$ effectively preserves the base-session knowledge, mitigating catastrophic forgetting.
5.3.4 Sensitivity Analysis of Hyper-parameters $\alpha$ and $\beta$
TABLE VI: Impact of hyper-parameters $\alpha$ and $\beta$ on performance in the miniImageNet dataset. “Average Acc.” represents the average accuracy across all sessions.
Impact of hyper-parameter $\alpha$
Value 40 80 120 140 160
Average Acc. 69.72 69.81 69.81 69.80 69.79
Impact of hyper-parameter $\beta$
Value 0.05 0.1 0.3 0.5 0.7
Average Acc. 69.64 69.70 69.71 69.81 69.79
We conducted a sensitivity analysis on the hyper-parameters $\alpha$ and $\beta$ of Eq. (22) in Table VI to evaluate their impact on our model’s average accuracy on miniImageNet. Each experiment varies one parameter with the other fixed. The results show that when $\alpha$ and $\beta$ are set within a proper range, the performance only varies slightly, demonstrating the model’s robustness to the hyper-parameters.
5.3.5 Impact of the Number of Training Samples
Figure 5: Performance across different sessions for 1-shot, 5-shot, 10-shot, and 20-shot learning scenarios on miniImageNet, CIFAR-100, and CUB-200 datasets.
We explored the effect of varying the number of training samples per class, or ”shots,” on Mamba-FSCIL’s performance. As shown in Fig. 5, increasing the number of training examples from 1-shot to 20-shot consistently improves accuracy across all sessions for the miniImageNet, CIFAR-100, and CUB-200 datasets. This demonstrates that Mamba-FSCIL benefits from the additional novel class training samples while effectively preserving base class knowledge, and preventing catastrophic forgetting.
5.3.6 Impact of the Number of Scan Paths
TABLE VII: Impact of the number of scan paths ($K$) on model performance for CIFAR-100 and CUB-200 datasets. “Base” refers to accuracy in the first session; “Final” to the accuracy in the last session; and “Avg” to the average accuracy across all sessions.
K CIFAR-100 CUB-200
[…]
1 82.35 55.94 67.23 80.41 60.84 67.48
2 82.65 57.13 67.53 80.49 61.08 67.86
4 82.80 57.51 68.14 80.90 61.65 68.13
We evaluated the influence of varying the number of scan paths ($K$) on Mamb-FSCIL’s performance on CIFAR-100 and CUB-200 datasets, as described in Table VII. Increasing the scan paths consistently enhances all performance metrics—base, final, and average session accuracies. Specifically, on CIFAR-100, elevating the scan paths from $1$ (top-left to bottom-right) to $2$ (adding bottom-right to top-left) boosts final session accuracy by 1.19%, from 55.94% to 57.13%, increases base session accuracy from 82.35% to 82.65%, and raises average accuracy from 67.23% to 67.53%. Extending to four paths (adding top-right to bottom-left and bottom-left to top-right) further improves final session accuracy to 57.51%, base session accuracy to 82.80%, and average accuracy to 68.14%. Observations on CUB-200 reflect similar trends. These findings highlight the substantial benefits of more scan paths in improving model performance, particularly their role in enhancing the integration of contextual information from multiple directions, which is crucial for handling complex visual tasks in dynamic learning environments. By default, we employed four scan paths in all experiments.
5.4 Comparison with Different Projection Architectures
TABLE VIII: Performance comparison of different projection architectures on CIFAR-100.
[…]
Dual Transformer 82.45 55.05 66.46
Dual MLP 82.57 56.02 67.57
Dual Selective SSM (ours) 82.80 57.24 68.08
To validate the effectiveness of the selective SSM projection branches in the Dual Selective SSM Projector for FSCIL, we performed a comparative experiment on the CIFAR-100 dataset. In this experiment, we replaced the selective SSM branches in our dual-branch architecture with MLP and Transformer structures while maintaining similar parameter counts across all models using a ResNet-12 backbone. The MLP setup includes three layers of multi-layer perceptrons, while the Transformer configuration incorporates a multi-head attention mechanism with 8 heads followed by a feed-forward network. Suppression loss with $\alpha=50$ is applied across all configurations and is directly applied to the output of the newly attached branch ($g^{\text{inc}}$) in the MLP and Transformer-based setups.
As shown in Table VIII, the Dual Selective SSM architecture consistently outperforms both MLP and Transformer setups across all critical metrics: base accuracy, final session accuracy, and average accuracy. Specifically, the Dual Selective SSM architecture enhances final session accuracy by 2.19% compared to the Transformer result and by 1.22% compared to the MLP setup. Furthermore, it achieves an average accuracy improvement of 1.62% over the Transformer and 0.51% over MLP.
5.5 Comparison with Dynamic Adaptation Method
Figure 6: Comparison of parameter growth (a) and session-wise accuracy (b) between Mamba-FSCIL and DSN [23].
We compared our Mamba-FSCIL with DSN [23] in Fig. 6, particularly in terms of parameter efficiency and session-wise performance. Both methods use the same ResNet18 backbone. As shown in Fig. 6 (a), Unlike DSN, which continually increases its parameter amount to learn new classes, Mamba-FSCIL armed with a dual selective SSM projector dynamically adjusts its parameters without expanding the parameter space. In addition, Fig. 6 (b) illustrates that Mamba-FSCIL consistently outperforms DSN in all sessions, demonstrating Mamba-FSCIL’s superior ability to retain old knowledge while learning new classes.
5.6 Visualization of Activation Maps
Figure 7: Activation maps for base and novel classes on the CUB-200 dataset. (a) Base classes; (b) Novel classes. Activation maps are generated using GradCAM from the $g^{\text{base}}$ and $g^{\text{inc}}$ branches.
We used GradCAM [118] to visualize the activation maps of a Swin Transformer-Tiny model trained on the CUB-200 dataset after 10 incremental sessions. The activation maps from the $g^{\text{base}}$ and $g^{\text{inc}}$ branches are visualized to assess their roles in handling base and novel classes.
Fig. 7 (a) shows that for images from the base classes, the $g^{\text{base}}$ branch significantly influences the final output, with activation regions effectively covering the main subject, indicating that this branch, fixed during incremental training, retains a strong representation of base classes. This confirms its role in preserving learned knowledge. In contrast, the $g^{\text{inc}}$ branch shows minimal activation, reducing interference with established knowledge. For novel classes, as shown in Fig. 7 (b), the fixed $g^{\text{base}}$ branch struggles to represent novel classes adequately, with incomplete activation coverage that could lead to classification errors. However, the $g^{\text{inc}}$ branch compensates by fully covering the subject, enhancing the model’s ability to represent novel classes. These visualizations demonstrate that the $g^{\text{base}}$ branch effectively preserves knowledge from base classes, while the $g^{\text{inc}}$ branch improves representation for novel classes without disrupting existing knowledge.
5.7 Analysis of Feature Embeddings
Figure 8: t-SNE visualization of feature embeddings from the CIFAR-100 dataset, showcasing base classes in (a) and (b) and novel classes in (c) and (d). Each color denotes a different class.
We analyzed feature embeddings using t-SNE visualizations for both base and novel classes from the test set of the CIFAR-100 dataset, comparing the strong baseline NC-FSCIL with our Mamba-FSCIL. Fig. 8 (a) and Fig. 8 (b) show the results of some base classes during the first and last sessions. It is shown that the base-class features of the baseline method are much scattered in the last session. As a comparison, Mamba-FSCIL demonstrates consistent intra-class compactness and distinct inter-class separation, indicating enhanced stability for base classes that better maintain their performance in incremental training. Fig. 8 (c) and Fig. 8 (d) depict the results of some novel classes in the 6-th and 8-th sessions. In both sessions, Mamba-FSCIL shows significantly more compact clusters with clearer margins among classes, highlighting its capacity to accommodate new classes. These results explain Mamba-FSCIL’s superior performance in the benchmarks.
6 Conclusion and Future Works
In this study, we propose the Mamba-FSCIL framework, leveraging selective state space models to address the challenges of FSCIL. Mamba-FSCIL utilizes a dual selective SSM projector to achieve dynamic adaptation without the need to continually expand the parameter space of the model. The dual design dynamically adjusts its parameters to integrate new classes while preserving the integrity of previously learned information. Furthermore, the incorporation of the class-sensitive selective scan mechanism, composed of the suppression loss and the separation loss, guides the dynamic adaptation process deliberately, enhancing both stability and adaptability in incremental sessions’ training. Empirical results demonstrate that Mamba-FSCIL outperforms the state-of-the-art static and dynamic methods across three benchmark datasets.
Future work for Mamba-FSCIL involves extending it to address highly correlated tasks like online class-incremental and cross-domain incremental learning. We also aim to adapt Mamba-FSCIL for multimodal scenarios, enhancing its effectiveness in dynamic and interactive environments. Additionally, exploring its application in embodied intelligence is also a key direction for future research.